### v1

# Adapting Multi-modal Large Language Model to Concept Drift in the Long-tailed Open World

Abstract
Real-world data often exhibit extreme imbalances and out-of-distribution (OOD) instances, which significantly biases the model training. While it has been extensively studied in vision and language domains separately, the impact of long-tailed open worlds on multi-modal large language models (MLLMs) has been largely overlooked. In this paper, we first demonstrate the susceptibility and vulnerability of vision-language models to significant biases caused by tail drift and out-of-distribution (OOD) drift during both the pre-training and fine-tuning stages. To eliminate the bias from different sources, we integrate the tailed drift adaptation and OOD drift detection into a unified framework by extending the concept drift theory to multi-modal. Specifically, a T-distribution-based drift adapter is proposed to effectively mitigate the bias induced by the long-tailed problem, which also facilitates the model in distinguishing OOD data through explicit distribution modelling. Extensive experiments show significant improvements in our model’s ability to adapt to tailed drift and OOD drift. Moreover, it enhances the efficiency and accuracy of image-text alignment in vision language model pre-training, particularly in the long-tail open world scenario. Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world scenario, to validate our findings. To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
1 Introduction
The explosive growth of data has precipitated a crisis for multi-modal large language models (MLLMs), which struggle to adapt to the long-tailed open world, encompassing the tailed drift and the out-of-distribution (OOD) drift. The tailed drift is brought by extremely imbalanced data, where the distribution of tailed categories shifts due to their inherent sparsity and noise, as well as the dominant head categories’ squeezing effect in the feature space. Meanwhile, OOD drift often occurs alongside the tailed drift and disrupts the underlying distribution of the model, as we continually encounter new concepts while navigating in an open world. Although the critical challenge of the long-tailed open world has been increasingly drawing attention in the realms of visual models [1] and language models [2], its impact on MLLMs, particularly vision-language (VL) models, has received scant attention. Therefore, we provide a comprehensive analysis of the impact of tail drift and OOD drift on VL models during both the pre-training and fine-tuning phases.
Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet [3] and the imbalanced dataset ImageNet-LT [1] is conducted. Due to the implicit feature centers of each category, we approximate them by averaging unit image and text features obtained by samples on the test set. To assess the intra-class compactness, the cosine distance between the image feature center and the text feature center from the same category is calculated and expressed as degrees. It is evident that training on the imbalanced dataset leads to a higher degree, indicating worse intra-class compactness brought by the tailed drift. Besides, with the tail drift intensifies, it results in a deterioration of the image-text alignment performance in tailed categories. Beyond the deterioration in tailed categories, the tailed drift also affects the image-text alignment in head categories with abundant training samples, which means that it leads to an overall performance degradation, not just the tailed categories. From the perspective of inter-class separability, we measure the average cosine distance from an image feature center to text feature centers of different categories. Figure 1(a) depicts that the VL model trained on the imbalanced dataset has lower inter-class separability. What’s more, we utilize KNN to extract 100 image and text feature centers of OOD samples to verify the impact of OOD drift on the pre-training of the VL model. Compared to training on the balanced dataset, the VL model trained under an imbalanced scenario is harder to distinguish between ID samples and OOD samples from the open world due to their similar inter-class separability. The undistinguished OOD drift will bias the underlying distribution of the feature space in the VL model, further disturbing the image-text alignment in the pre-training.
(a) Image-Text Alignment in Pre-training
(b) Feature Space Allocation in Fine-tuning
Figure 1: The impacts of tailed drift and OOD drift on the vision language model in the stages of pre-training and fine-tuning, respectively. (a) In terms of the pre-training, we visualize the alignment results on both a balanced dataset (denoted as BL) and an imbalanced dataset (as LT). The cosine metric is used to measure the distances between unit image and text features across various categories including OOD samples, which is expressed as degrees. A smaller degree indicates a higher level of similarity between the features. Thus, it provides a feature-level visualization of the intra-class compactness and inter-class separability in the vision language model. (b) In the context of fine-tuning, the mutual cosine distance between the centers of each category in the classifier is directly visualized to illustrate the feature space of the classifier, denoted as blue bars. Besides, the average cosine distance between each category center and OOD samples is calculated, which is represented as orange bars.
Fine-tuning: We explicitly leverage the weights of the embedding layer in the VL model to visualize its feature space. The average cosine distance between each category and others is calculated as exhibited in the blue bars of Figure 1(b). with the decreasing of the training samples, a smaller degree means a worse inter-class separability. It is verified that tail drift leads to a compression of the feature space for tailed categories with a limited number of training samples, while head categories with abundant samples dominate the overall feature space of the VL model. Moreover, the average cosine distance between each category and unit features extracted by OOD samples is applied to reveal the OOD separability as denoted as orange bars in Figure 1(b). Since head categories occupies most of the feature space, OOD samples are closer to the center of the head categories compared to the tail categories, implying that in the stage of fine-tuning, it is difficult for the VL model to distinguish between OOD samples under imbalanced scenarios.
To effectively address tailed drift and OOD drift within a unified framework, which often occur simultaneously, we encapsulate them using the concept drift theory. Therefore, summarizing the above challenges of vision language models in the long-tailed open world, it raises the important question:
How to adapt multi-modal large language model to concept drift in the long-tailed open world?
Remark 1.1.
Research Objective: Our focus lies in addressing the drift that MLLMs exhibit when confronted with the long-tailed open world, rather than leveraging MLLMs to enhance classification performance on long-tailed open datasets. The classification is exploited as the downstream task to provide a more intuitive visualization of the concept drift suffered by the MLLMs, which could have other downstream tasks.
Therefore, we propose a concept drift-aware multi-modal large language model, mitigating the tail drift and OOD drift encountered in the long-tailed open world. Firstly, we introduce the concept drift theory to the multi-modal domain, which provides a more holistic perspective to explain tailed drift and OOD drift. Then, the T-distributed adapter is proposed to be embedded in the hyperspherical feature space. It aligns image-text features for contrastive learning in pre-training. The desirable light-tailed property of the proposed T-distributed spherical metric (tSP) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift. Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution. The proposed tSP distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings. And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the tSP distribution. Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
In summary, our paper mainly makes the following contributions:
  1. 1.
    We are the pioneers in revealing the unexplored impacts of long-tailed open world to multi-modal large language models, especially in the image-text alignment in the pre-training and feature space allocation in the fine-tuning. This allows future research to more comprehensively study the impact of defect data on MLLMs.
  2. 2.
    The concept drift theory is introduced and extended to multi-modal, integrating the tailed drift and OOD drift in a unified framework. And the T-distributed spherical adapter is proposed to perform the tailed adaptation and OOD detection in the pre-training and fine-tuning stage of the VL model.
  3. 3.
    Extensive experiments evaluate the performance of our method under the long-tailed open world. Compared to specialized models, ours demonstrates superior performance in downstream tasks of long-tailed classification and OOD detection. Crucially, our model effectively addresses drift in image-text alignment, facilitating large-scale pre-training of MLLMs.
  4. 4.
    We build a group of multi-modal datasets OpenMMlo under the long-tailed open world, by extending existing image-based long-tailed open datasets. It contains about 740k image-caption pairs with related category annotations. To support and encourage the community focused on multi-modal, we have made both the OpenMMlo and our code public.
2 Methodology
2.1 Multi-modal Concept Drift Theory
Concept drift is a phenomenon in which the statistical properties of a target domain change over time in an arbitrary way [4]. Formally, given a set of examples denoted as the data stream $S_{0,t}=\{d_{0},...,d_{t}\}$, where $d_{i}=(X_{i},y_{i})$ is one data instance, $X_{i}$ and $y_{i}$ respectively denote the feature vector and the label, and $t$ represents the timestamp of the instance in the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(X,y)$. The concept drift is formalized as: $\exists t:P_{t}(X,y)\neq P_{t+1}(X,y)$, where the joint probability $P_{t}(X,y)$ can be decomposed as $P_{t}(X,y)=P_{t}(X)\times P_{t}(y|X)$. Although the concept drift due to tailed and OOD data often co-occur, they are fundamentally distinct phenomena. The tailed concept drift foucus on the drift in $P_{t}(X)$, while $P_{t}(y|X)=P_{t+1}(y|X)$ remains unchanged. However, the OOD drift from the unknown categories triggers the drift of both $P_{t}(y|X)$ and $P_{t}(X)$, that $P_{t}(y|X)\neq P_{t+1}(y|X)$ and $P_{t}(X)\neq P_{t+1}(X)$. Therefore, concept drift theory provides a unified framework to harmonize the tailed shift and OOD shift that often occur together, enabling more robust and adaptive deep learning models.
In the context of multi-modal vision language models, we extend the concept drift theory from a single data stream to multiple data streams. Each data modal is associated with a distinct data stream. Thereby, the multi-modal concept drift framework can robustly handle the complex, heterogeneous data distributions inherent to vision language models. Therefore, we formally define multi-modal concept drift as follows:
Definition 2.1.
Assume that there are $N$ data streams corresponding to $N$ modalities, given a set of examples denoted as $S_{0,t}=\{S_{0},...,S_{i},...,S_{t}\}$, where $S_{i}=(s_{1},...,s_{j},...,s_{N})$ and $s_{j}=(X_{ij},y_{i})$ is one data instance from a single $j$-th data stream, $X_{ij}$ is the feature vector, $y_{i}$ is the label and $t$ is the timestamp of the data stream. $S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $F_{0,t}(S_{i})\neq F_{t+1,\infty}(S_{i})$, denoted as:
$\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i})$ (1)
Remark 2.1.
Differences: Concept Drift vs. Data Drift (Covariate Drift) Data drift entails changes solely in the distribution of inputs $P(x)$, while concept drift involves alterations in both input and output distributions, i.e., $P(x)$ and $P(y)$, leading to changes in the decision boundary. Furthermore, data drift predominantly stems from internal factors like data collection and processing, whereas concept drift typically arises from external factors, reflecting real-world changes.
2.2 T-distributed Adapter for Long-Tailed Open World
Within the concept drift framework, data distribution-based drift detection and adaptation have emerged as the dominant approaches [5, 6]. To effectively mitigate the concept drift in the long-tail open world, we integrate directional statistics into distribution-based drift detection and adaptation, and propose a T-distributed adapter to alleviate it in the vision language model.
2.2.1 Directional statistics
Directional statistics primarily focus on the distribution of eigenvector angles, while neglecting the impact of eigenvector module lengths. Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set. A key idea in directional distribution is the tangent-normal decomposition. Any unit vector $x$ can be decomposed as:
$x=t\mu+(1-t^{2})^{\frac{1}{2}}v,t\in[-1,1]$ (2)
with $v\in\mathbb{S}^{d-2}$ a tangent to $\mathbb{S}^{d-1}$ at $\mu$ [7, 8], where $v$ and $t$ are independent and $v$ is uniform on $\mathbb{S}^{d-2}$. Thus, the intersection of $\mathbb{S}^{d-1}$ with the hyperplane through $t\mu$ and normal to $\mu$ is a $(d-2)$-dimensional sphere of radius $\sqrt{1-t^{2}}$, that $t$ has density as following:
$p_{T}(t;d)\propto(1-t^{2})^{\frac{d-3}{2}},t\in[-1,1]$ (3)
Therefore, through the marginal density $p_{T}$ and $p_{v}$, we can estimate the density of the entire spherical distribution. One prominent instance is the von Mises-Fisher distribution (vMF) [9], which can be interpreted as a probability distribution over the cosine similarity between a unit vector $x$ and a fixed mean direction $\mu$, following the density:
$p_{X}(x;\mu,\kappa)\propto\exp{(\kappa\mu^{T}x)}$ (4)
where $\kappa\geqslant 0$ denotes the concentration and $\exp$ represents the exponential function. Therefore, combined with the Eq. 2 and Eq. 3, the density of vMF is:
$p(x)=C_{X}(\kappa,d)^{-1}\exp{(\kappa\mu^{T}x)},\quad x\sim\text{vMF}(\mu,\kappa)$ (5)
$C_{X}(\kappa,d)=\frac{(2\pi)^{d/2}I_{d/2-1}(\kappa)}{\kappa^{d/2-1}}$
where $I_{m}$ denotes the modified Bessel function of the first kind at order $m$.
2.2.2 T-distributed Adapter
Although the spherical nature of the vMF mitigates the bias induced by module length in long-tail open worlds, its exponential property, as evident in Eq. 4, leads to a winner-take-all behaviour, where only a marginally higher logit than its competitors will dominate the eigenspace, receiving an exponentially disproportionate share of the eigenspace. As a consequence, the head categories, which possess abundant samples, exert a dominant influence over the entire model and squeeze the eigenspace of tailed categories, causing the crowding problem.
Therefore, inspired by the T-SNE [10], we design a T-distribution based spherical metric (tSP), which follows the density:
$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)}$ (6)
where the unit feature vector $x\in\mathbb{S}^{d-1}$ and the center $\mu\in\mathbb{S}^{d-1}$. Through Eq. 3, we can get the marginal normalizer:
$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (7)
where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. 2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
$N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$ (8)
Thus, the probability density function of the proposed tSP distribution is as follows:
$p(x)=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)},\quad x\sim\text{tSP}(\mu)$ (9)
The detailed derivation process is provided in Appendix A.2.
Figure 2: The proposed T-distributed spherical metric with various $\kappa$ and the classical vMF metric when $\kappa=1$.
In terms of adapting to tailed concept drift, the tSP metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 2. The high kurtosis of tSP is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers. Formally, $L_{\text{tsp}}(\mu,x)=\frac{2}{\kappa(1-\mu^{T}x)+\epsilon}$ denotes the T-distributed spherical metric where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric. Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$. Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{tsp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$. Consequently, the light-tailed tSP distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
Likewise, the tSP metric is directly applied to detect the OOD concept drift. A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace. Following [11], a simple non-parametric KNN is adopted to partition the data into two sets (ID vs. OOD), which does not impose any distributional assumption on the feature space. Here the distance is the tSP metric with respect to the k-th nearest neighbor.
Figure 3: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks. By framing multi-modal heterogeneous data as concept drift data streams, we unify both tail drift and OOD drift. In the pre-training, the tSP metric aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin. Coupled with the language model loss, they drive the training of all the module. In the downstream task, the image encoder and the text decoder are frozen out of training, and two linear layers as the projector embed features into the hyperspherical eigenspace. The T-distributed adapter explicitly incorporates light-tail property into the feature space modelling, allowing it to effectively adapt tail drift, and performs OOD drift detection based on the distribution.
2.3 T-distributed Vision Language Model for the Long-Tailed Open World
As illustrated in Figure 3, our proposed vision language model contains two stages, the pre-training and the fine-tuning on the downstream task. Specifically, the classification task is chosen to visualize the impact of the bias caused by the long-tailed open world on the model. The multi-modal concept drift theory offers a unified framework for integrating tail drift adaptation and OOD drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
The architecture of the proposed vision language model contains the image encoder, text encoder and text decoder. With an input image $I_{i}\in\mathbb{R}^{H\times W}$, the visual features $x_{\text{img}}$ are extracted by the image encoder $E_{\text{img}}$ and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
$x_{\text{img}}=P_{\text{norm}}(E_{\text{img}}(I_{i}))\in\mathbb{R}^{n\times d}$ (10)
where $n$ is the number of visual features and $d$ represents the feature dimension. The image encoder $E_{\text{img}}$ can be any common visual backbones, such as Vit-Base [12], Vit-Large [12] and ResNeXt-50 [13]. In terms of the text encoder $E_{\text{txt}}$, with a processed input text sequence $T_{i}$, text features are extracted by the language encoder and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
$x_{\text{txt}}=P_{\text{norm}}(E_{\text{txt}}(T_{i}))\in\mathbb{R}^{n\times d}$ (11)
where $n$ is the number of input tokens and $d$ represents the feature dimension. In our case, Bert [14] is used as the language encoder.
In the pre-training of the vision language model, the T-distributed adapter aligns image and text features by contrastive learning. It seeks to align visual and textual transformer feature spaces by promoting similar representations for positive pairs and dissimilar representations for negative pairs. More importantly, our approach circumvents model bias stemming from the long-tailed distribution of data. It couples with the language modelling loss to drives the training of all three modules.
In the downstream classification task, we leverage the T-distributed adapter to mitigate the long-tailed distribution of categories, thereby improving classification performance. Furthermore, by explicitly modelling the feature space, the T-distributed adapter enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples. Besides, following the Blip [15], the image encoder and the text decoder are frozen out of training during fine-tuning. A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
2.4 Building Multi-modal Datasets Group OpenMMlo for the Long-Tailed Open World
Due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to follow the long-tailed distribution in the open world. Recognizing the demand for higher-quality multi-modal data with long-tailed distribution in an open world, we developed a group of datasets called Open Multi-modal Long-Tailed OOD Datasets (OpenMMlo).
We extend the open-source datasets, namely ImageNet-LT [16], iNatualist2018 [17] and Places-LT [16]. ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category. Besides, it consists of 18k images for OOD detection. Places-LT has 184.5K samples from 365 classes, with class samples ranging from 4,980 to 5. The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes. We use the InstructBLIP [18] to generate the related caption of the image, with the prompt of "What does this picture describe? Please describe in detail its size, location, color, and its relationship to the surroundings.". For more details about OpenMMlo, please refer to Appendix A.4.
3 Experiments
In this section, we first present the performance in downstream long-tailed classification and OOD detection tasks, which is induced by tail drift and OOD drift on MLLMs. Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift. The constructed long-tailed multi-modal datasets OpenMMlo is utilized for training and validating. In terms of the OOD drift detection, we follow the setting of CIDER [19]. The model is trained on CIFAR100-LT [20] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [21], Places365 [22], LSUN [23], iSUN [24] and Texture [25]. More detailed experimental implementations are given in Appendix A.3.
3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
We compare our proposed vision-language model with other specialized models to explicitly demonstrate its superior performance in long-tailed open-world scenarios. As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across all three large-scale datasets. The results validate the robustness of our vision language model against biases arising from tail drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
Table 1: Evaluation results of long-tailed classification on ImageNet-LT, iNatualist2018 and Places-LT datasets with Vit-Large. The best-performing models are highlighted in red. Many, Medi. and Few denote the evaluated splits of many-shot (>100 training samples), medium-shot (20-100 samples) and few-shot (<20 samples). Top-1 accuracy is applied to evaluate the performance of different methods.
Methods ImageNet-LT iNaturalist 2018 Places-LT
Many Medi. Few All Many Medi. Few All Many Medi. Few All
MiSLAS [26] 62.0 49.1 32.8 51.4 - - - 70.7 - - - 38.3
Bl-Softmax [27] 64.1 48.2 33.4 52.3 - - - 70.6 - - - 39.4
LADE [28] 64.4 47.7 34.3 52.3 - - - 69.3 - - - 39.2
ACE [29] - - - 56.6 - - - 72.9 - - - -
RIDE [30] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6 - - - -
PaCo [31] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2 36.1 47.9 35.3 41.2
NCL [32] - - - 60.5 - - - 74.9 - - - 41.8
MEM [33] 76.0 62.0 43.7 64.9 83.6 83.0 81.3 82.4 49.1 47.4 37.4 46.0
Ours 80.3 71.1 37.2 65.6 84.3 83.1 80.7 82.6 61.6 53.3 32.6 46.9
Furthermore, our model exhibits superior performance on the Many and Medium splits, while it slightly lags behind other models in the Few split. On one hand, our findings align with the scaling law observed in large models, where the performance exhibits a power law relationship with the size of the data. As the amount of data increases, the model’s feature extraction capability becomes more robust leading to improved performance. On the other hand, our goal is to alleviate the bias induced by tail drift, instead of averaging the feature space of each category and sacrificing head categories to enhance the performance of tail categories.
[…]
Methods SVHN LSUN iSUN Texture Average
FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC
ProxyAnchor [34] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0 65.0 86.0
CE + SimCLR [35] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0 52.9 87.3
CSI [36] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5 64.6 87.0
SSD+ [37] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2 64.5 87.4
KNN+ [11] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4 55.1 88.3
CIDER [19] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3 31.1 93.0
Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3 26.6 96.9
In terms of OOD drift detection, our proposed vision language model, trained on CIFAR100-LT as an in-distribution dataset, demonstrates exceptional performance across four diverse OOD datasets, as shown in Table 2. Our approach stands out with two significant advancements. Firstly, the training of our model does not incorporate any additional data from the open world to delineate the decision boundary between ID samples and OOD samples. Secondly, our proposed model detects OOD drift based on the hyperspherical distribution, without the need for any specialized modules. The proposed methodology offers the convenience of training and inference for large models.
Table 3: Evaluation results of image-text alignment of different contrastive learning strategies in the stage of pre-training, from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories. The cosine metric is utilized to measure these distances, which is expressed as average degrees with standard deviation in brackets. We compare our proposed tSP with classical cosine loss, under balanced scenario (BL, ImageNet) and imbalanced scenarios (LT, ImageNet-LT), respectively.
Pre-training
ID Intra-class
Compactness $\downarrow$
ID Inter-class
Separability $\uparrow$
ID vs. OOD
Separability $\uparrow$
BL / Cosine 33.0 ($\pm$2.85) 84.3 ($\pm$1.64) 98.4 ($\pm$0.98)
LT / Cosine 49.2 ($\pm$4.95) 76.5 ($\pm$2.16) 80.7 ($\pm$1.90)
LT / tSP 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training. As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/tSP. It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios. Notably, the decrease in standard deviation demonstrates that the model considerably mitigates the bias induced by tail drift. In addition, our method achieves remarkable inter-class separability within in-distribution categories under long-tailed scenarios, even surpassing the performance of cosine achieved under the balanced dataset. It confirms the effectiveness of the proposed high kurtosis method in enhancing the alignment between images and text in the pre-training stage. Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter. It ensures our approach performs robust for OOD drift detection based on the spherical distribution robustly.
3.3 Ablation Experiments
3.3.1 T-distributed Spherical Embedding in the Pre-training and Fine-tuning
Table 4: Ablation evaluation results with or without the T-distributed adapter in the pre-training or fine-tuning. The ✓denotes the stage is trained with the T-distributed adapter. The results are based on the ImageNet-LT with the Vit-base. Top-1 accuracy (Acc) is used as the metric.
Pre-training Fine-tuning Acc
- - 53.5
- 55.7
- 60.0
63.8
Firstly, we conduct ablation experiments to verify the improved performance of the T-distributed adapter in the stage of pre-training and fine-tuning, respectively. As demonstrated in Table 4, our proposed T-distributed adapter exhibits improvements in both the pre-training and fine-tuning stages of the vision language model. It is worth highlighting that the T-distribution adapter plays a more prominent role during the fine-tuning stage. We argue that it is due to different characteristics of the pre-training and fine-tuning. During fine-tuning, the model is directly involved in specific downstream tasks, and explicit category centers are present in the classifier. In contrast, pre-training primarily focuses on aligning image-text features, where the implicit information of categories is embedded. As a result, the T-distribution adapter’s impact is more pronounced in the fine-tuning stage compared to pre-training.
3.3.2 Various Concentration Parameter $\kappa$
Furthermore, we conduct ablation experiments to examine the impact of different concentrations of the T-distributed adapter on the overall performance of the vision language model in Table 5. Four fixed degrees of concentration are involved, namely $\kappa=4$ 16, 64 and 128. The greater the degree of concentration, the greater the kurtosis of the T-distributed spherical metric. Besides, the concentration can also be utilized as a trainable parameter joining in the training, with the initial setting of 16.
Table 5: Ablation evaluation results of various concentration parameter $\kappa$ on long-tailed classification task. "Training" denotes the $\kappa$ involved in the training as a parameter of the model with an initial setting of 16. The results are based on the ImageNet-LT with the Vit-base. Top-1 accuracy (Acc) is used as the metric.
$\kappa$ 4 16 64 128 Training
Acc 61.3 63.8 59.9 57.2 62.3
In Table 5, setting the concentration parameter to $\kappa=16$ yields superior results. We argue that, a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model. In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the tSP metric. In the context of concentration as a trainable parameter with the initialization of 16, there is a slight reduction in model performance, accompanied by a marginal increase of the concentration parameter to 16.37. We assert that, the introduction of the new parameter increases the model’s complexity, thereby making the training process more challenging. Additionally, the low learning rate hinders the effective training of the concentration parameters, leading its less change compared with the initialization.
4 Conclusions And Outlook
Our findings indicate that visual-language models are significantly affected by biases introduced during both pre-training and fine-tuning in long-tailed open-world scenarios. To address this, we propose a concept drift-aware unified framework for visual-language models. This framework incorporates a T-distributed adapter designed to mitigate biases arising from both tailed drift and out-of-distribution (OOD) drift. Additionally, we introduce a comprehensive set of multi-modal datasets (OpenMMlo) tailored to the long-tailed open world, which includes images, captions and related category annotations.
We hope that our work will inspire future advancements in multi-modal large language models, specifically addressing the mitigation of biases originating from real-world data challenges, such as tailed drift and OOD drift. In future research, we will leverage concept drift theory to analyze the temporal, intensity, and regional aspects of the various drifts.
Appendix A Appendix / supplemental material
A.1 Related Works
A.1.1 Multi-modal Large Language Model
Large Language Models (LLMs) have recently significantly impacted the field of natural language processing. Through alignment techniques such as supervised learning and reinforcement learning with human feedback, LLMs can effectively generalize to perform a wide range of tasks, even with limited training data. A remarkable application of LLM is ChatGPT, which presents an amazing ability to interact with humans. OpenAI’s ChatGPT and GPT4 are prime examples of the impact that AI can have, and there have been extensive open-source efforts to replicate their success, such as OPT [38], BLOOM [39], PALM [40], LLaMA [41].
Multi-modal large language models have further promoted the development of the vision-language model [42, 43, 44, 45, 46, 47, 48, 49]. CLIP [42] was introduced to separately extract features from different encoders and combine them using contrastive learning. Building on CLIP, GLIP [43] was developed to learn object-level, language-aware, and semantic-rich visual representations, unifying object detection and phrase grounding for pre-training. Different from the contrastive method, Flamingo [44] aligned a pre-trained vision encoder and language model using gated cross-attention, demonstrated impressive few-shot learning capabilities. BLIP2 [45] was subsequently introduced, and it employed a Flan-T5 [50] along with a Q-Former to effectively align visual features with the language model. MiniGPT4 [46], the most recent development in the field is the PaLM-E model, which features 562 billion parameters and is designed to integrate real-world continuous sensor modalities into an LLM, thereby establishing a connection between real-world perceptions and human languages. Based on Visual Fundamental Models like BLIP mentioned above, Visual ChatGPT adopts ChatGPT as the central component for interacting with users. It integrates multiple visual foundation models and utilizes prompt engineering, also known as Prompt Manager, to instruct ChatGPT about the usage, input-output format, and capabilities of each foundation model. This enables ChatGPT to determine how to invoke these models to fulfill the user’s requirements. Besides, GPT-4V(ision)[51] and GPT-4O(mni) have recently shown unprecedented ability in understanding and processing an arbitrary mix of input images and texts.
A.1.2 Long-tailed Open World
In vision tasks, significant efforts have been devoted to mitigating the challenges posed by the long-tailed open world. Two prominent research directions have emerged: long-tailed classification under open-world settings, exemplified by approaches like OLTR++ [16, 1], LUNA [52], DALC [53], Open-sampling [54] and TLC [55], and OOD detection in long-tailed recognition, as seen in methods such as PASCL [56], EAT [57]. OLTR++ [16, 1] proposed an ensemble algorithm, consisting of dynamic meta-embedding to improve the recognition of tail categories and active learning for open categories detection. LUNA [52] presented a distribution-sensitive loss to weigh more on the tail classes and a local-density-based metric to measure the novelty of OOD samples. DALC [53] designed an active distribution optimization algorithm for clustering, querying and classification to balance the classification bias. Open-sampling [54] rebalances class priors by sampling labels from a complementary distribution for each open-set instance, mitigating class imbalance. TLC [55] utilizes Dempster-Shafer Evidence Theory in a multi-expert framework for uncertainty estimation of tail and OOD samples. PASCL [56] applied supervised contrastive learning to explicitly boost the model to distinguish between tail-class in-distribution samples and OOD samples. EAT [57] introduces abstention classes for clear decision boundaries and augmenting tail classes with context-rich OOD data to focus on discriminative features. MCM [58] pioneers the integration of vision language models into OOD detection, enabling zero-shot OOD by aligning visual features with text concepts through a proposed maximum concept matching approach.
Additionally, in the domain of the language model, [2] corroborates that large language models (LLMs) also struggle to learn long-tailed knowledge. While larger models are better at absorbing the long-tailed knowledge, they estimate that current models must be scaled by many orders of magnitude to reach competitive performance. Besides, [59] alleviates the long-tail problem in neural machine translation by quantifying token classification and sequence generation, and introduces an anti-focus loss that incorporates beam search inductive biases to better adapt model training to conditional text generation.
A.1.3 Concept Drift
In the review [4], the algorithms related to concept drift are categorized into three groups: error rate-based, data distribution-based and multiple hypothesis-based. Our proposed algorithm belongs to the distribution-based concept drift detection and adaptation method. Distribution-based concept drift algorithms not only accurately detect drift through explicit distributions but also analyze the drift to identify its happening timing, location, and severity.
Besides, RBM-IM [60] proposes a novel trainable concept drift detector based on Restricted Boltzmann Machine, to slove the concept drift in multi-class imbalanced data streams. CALMID [61] propose a comprehensive active learning method for multiclass imbalanced streaming data with concept drift, including an ensemble classifier, a drift detector, and a variable threshold uncertainty strategy. DES-ICD [62] is a dynamic ensemble selection method for imbalanced data streams with concept drift. It considers local performances of base classifiers and addresses class imbalance using a novel synthetic minority oversampling technique. GOOD [63] develops an graph OOD benchmark, which explicitly distinguishes between covariate and concept shifts and design data splits that accurately capture these different shifts.
A.1.4 Hyperspherical Distribution Modelling
The Bayesian estimation of vMF mixture model with variational inference is addressed in [64]. The learning task in VI consists of the optimization of the variational posterior distribution. Besides, a deep metric learning model for image classification and retrieval is presented in [65] , which utilizes the vMF distribution to define the loss function and introduces an effective alternative learning algorithm by updating class centers. The model captures global information in the embedding space and approximates the class distribution during training, leading to improved performance in image tasks. [66] extends the vMF distribution to regularize the intra-class feature distribution for imbalanced, small-scale and noisy data. [67] focus on using hyperspherical embedding to alleviate the crowding problem arisen by the imbalanced data. [19] utilizes hyperspherical embeddings for OOD detection in representation learning, consisting of two losses, a dispersion loss to increase angular distances between different class prototypes, and a compactness loss to ensure samples are closer to their respective class prototypes.
A.2 The T-distributed Spherical Distribution
Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set. The proposed T-distribution based spherical metric (tSP) follows the density:
$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)}$ (12)
where $x\in\mathbb{S}^{d-1}$, direction $\mu in\mathbb{S}^{d-1}$ and concentration $\kappa\in\mathbb{R}_{\geq 0}$. Let $T$ bet a random variable that denotes the dot-product $t=\mu^{T}x$, then $T=2Z-1$, with $Z\sim\text{Beta}(\alpha,\beta)$, where $\alpha=\frac{d-1}{2}$ and $\frac{d-3}{2}$.
Proof.
Given Eq. 3, the marginal distribution of the dot-product $t$ is
$t\propto\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}$ (13)
So, its normalizer is:
$N_{T}(\kappa,d)$ $=\int_{\mathbb{S}^{d-1}}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}\mathrm{d}t$ (14)
$=\int_{-1}^{1}\frac{1}{\kappa(1-t)}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-3}{2}}\mathrm{d}t$
$=\frac{1}{\kappa}\int_{-1}^{1}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}\mathrm{d}t$
Given the useful integral function:
$\int(1+x)^{a}(1-x)^{b}\mathrm{d}x=2^{a+b+1}B_{\frac{x+1}{2}}(a+1,b+1)+C$ (15)
So, its normalizer is:
$N_{T}(\kappa,d)$ $=\frac{1}{\kappa}2^{d-3}(B_{1}(\frac{d-1}{2},\frac{d-3}{2})-B_{0}(\frac{d-1}{2},\frac{d-3}{2}))$ (16)
$=\frac{1}{\kappa}2^{d-3}B(\frac{d-1}{2},\frac{d-3}{2})$
The Beta function:
$B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$ (17)
So, the normalizer is
$N_{T}(\kappa,d)=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (18)
where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. It follows that the probability density function of the marginal distribution of the dot product is
$p_{T}(t;\kappa,d)$ $=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}$ (19)
$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}$
$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(2z)^{\frac{d-1}{2}-1}(2-2z)^{\frac{d-3}{2}-1}$
$=\frac{2}{\kappa}B(\alpha,\beta)^{-1}z^{\alpha-1}(1-z)^{\beta-1}$
where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. ∎
Due to the surface area of the hyper-sphere $\mathbb{S}^{d-1}$ is:
$A_{d-1}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}$ (20)
The T-distributed sperical distribution is expressed via the tangent normal decomposition as a joint distribution between $T\sim p_{T}{t;\kappa,d}$ and $V\sim\mathcal{U}(\mathbb{S}^{d-2})$. Since $T\perp\!\!\!\perp V$, the tSP normalizer $N_{x}(p,k)$ is the product of the normalizer of $p_{T}(t;\kappa,d)$ and the uniform distribution on $\mathbb{S}^{d-2}$ is:
$N_{X}(\kappa,d)$ $=N_{T}(\kappa,d)\cdot A_{d-2}$ (21)
$=2^{\alpha+\beta-1}B(\alpha,\beta)\frac{2\pi^{\beta}}{\kappa\Gamma(\beta)}$
$=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$
Thus,
$p_{X}(x;\mu,\kappa)=N_{X}(\kappa,d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)}$ (22)
A.3 Implementation Details
For our language-guided image tokenizer, we leverage the strengths of both BERT [68] and ViT as our text encoder, text decoder and visual encoder, respectively.
We employ ViT-Large as our visual encoder, which consists of 24 transformer encoder layers and an FFN intermediate size of 4,096. The input image size is set to $384\times 384$, with a patch size of $16\times 16$. The hidden dimensions of the ViT-Large are 1,024, with 16 attention heads. And, the number of parameters is about 307 million. Besides, we also use Vit-Base and ResNeXt-50 to perform ablation experiments. Vit-Base consists of 12 transformer encoder layers and an FFN intermediate size of 3,072. The hidden dimensions of the ViT-Base are 768, with 12 attention heads. The number of parameters is about 86 million. The input image size is set to $224\times 224$. In addition, ResNeXt-50 has 16 residual blocks with 50 layers. Each block has 3 convolutional layers with the kernel size of $3\times 3$, the stride of 1 and the padding of 1. The batch normalization and max pooling are utilized to connect the convolutional layers. The classification head hidden dimensions are 2,048.
Additionally, BERT as the language model in our vision-language model, has 12 transformer layers with 768 hidden dimensions and 3,078 intermediate dimension. The number of attention heads is 12, with the input sequence length of 512. It has approximately 110 million parameters.
(a) Sample in Training Set
(b) Sample in Test Set
(c) Sample in Open Set
Figure 4: Samples of OpenMMlo in training set, test set and open set.
In terms of the pre-training progress, the hyperparameters are presented in Table 6. We utilize the AdamW optimizer, which is configured with a cosine annealing schedule as the learning policy. The initial learning rate is set to $2\times 10^{-5}$, and the AdamW optimizer is employed with hyperparameters $\beta=(0.9,0.98)$. Additionally, we set the weight decay to 0.05 and the dropout rate to 0.1. During the first 1,000 warm-up steps, the learning rate increases to $2\times 10^{-5}$, and subsequently decays to $10^{-7}$. Unless otherwise specified, the pre-training of our vision language model consists of 800,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs, which takes approximately 7 days to complete.
Table 6: The training hyperparameters of our vision language model.
Pre-training
Training Steps 400,000
Warmup Steps 1,000
Optimizer AdamW
Learning Rate 1e-4
Learning Rate Decay Cosine
Adam $\beta$ (0.9, 0.98)
Weight Decay 0.05
Batch Size 50
Fine-tuning
Training Steps 18,000
Warmup Steps 0
Optimizer AdamW
Learning Rate 2e-5
Learning Rate Decay Cosine
Adam $\beta$ (0.9, 0.98)
Weight Decay 0.05
Batch Size 400
While in the fine-tuning on downstream task of classification, the initial learning rate is reduced to $10^{-6}$ without the warmup. The visual encoder and text decoder are frozen out of the training. Thus, the batch size can be increased to 400. The fine-tuning consists of 18,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs, which takes approximately 15 hours to complete. Other training parameters are as the same as the pre-training.
When evaluating the performance of our VL model under the long-tailed open world, we use the top-1 accuracy metric on the downstream classification task. In particular, the categories are split into three groups: many-shot (with more than 100 training samples), medium-shot (with 20-100 training samples), and few-shot (with fewer than 20 training samples). The Top-1 accuracies are computed for each group to evaluate the performance of mitigating the bias introduced by the long-tail distribution, respectively. Furthermore, in order to assess the capability of detecting the OOD drift, we employ two metrics: FPR95 which measures the false positive rate of OOD samples when the true positive rate of ID samples reaches 95%, and AUROC providing the area under the receiver operating characteristic curve. Besides, cosine distance is exploited to measure the distances between features and centers in the feature space of the VL model.
A.4 Building Multi-modal Long-Tailed OOD Datasets Group OpenMMlo
Figure 4 showcases the samples utilized for training and validation in our study. To intuitively verify the impact of long-tail open world scenarios on multi-modal large language models, we employ classification as our downstream task. When matching images and texts, we strategically mask words that are directly related to category names. This approach ensures the accuracy and reliability of our experimental results. As depicted in Figure 4, comprehensive descriptions of the image are provided through long-form text, encompassing details such as size, position, color, relationships, and other relevant information about the objects present in the image. This ensures a detailed and information-rich depiction of the visual content.
We have publicly released the datasets used for training and validation, as well as the original unmasked datasets.
A.5 More Experiment Results
A.5.1 Various Image Encoder
Table 7: Ablation evaluation results of various image encoders, including the ResNeXt50, Vit-base and Vit-Large. The results are based on the ImageNet-LT. Top-1 accuracy is used as the metric.
Methods ResNeXt-50 Vit-Base Vit-Large
CE 45.2 53.5 59.7
$\tau$-normalized [69] 49.0 57.1 62.7
cRT [69] 49.3 58.1 64.0
LWS [69] 48.2 57.9 63.3
MEM [69] 49.6 58.9 64.9
Ours 52.2 63.8 65.6
Beyond leveraging Vit-Large as the backbone, we also explore alternative image encoders in ablation experiments, namely Vit-Base, Vit-large and ResNeXt-50 to further dissect the model’s performance. We conducted ablation experiments on various image encoders to demonstrate the effectiveness of our method across different backbone architectures, as illustrated in Table 7. We select various backbones consisting of ResNeXt50, Vit-base and Vit-large, with different model size. The results confirm that our method consistently achieves superior performance across various backbone networks.
A.5.2 More Results
In this section, we provide more results of OOD detection.
[…]
Methods SVHN LSUN iSUN Texture Average
FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC
MSP [70] 78.9 79.8 83.5 75.3 84.6 74.5 86.5 72.5 84.4 74.2
ODIN [71] 70.2 84.9 76.4 80.1 79.5 79.2 85.3 75.2 82.2 75.2
Mahalanobis [72] 87.1 80.6 84.2 79.4 83.2 78.8 61.7 84.9 84.6 73.9
Energy [73] 66.9 85.3 59.8 86.7 66.5 84.5 79.0 80.0 81.4 76.4
GODIN [74] 74.6 84.0 93.3 67.2 94.3 65.3 86.5 69.4 89.1 69.0
LogitNorm [75] 59.6 90.7 81.1 83.0 84.2 80.8 86.6 75.6 80.3 78.6
ProxyAnchor [34] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0 65.0 86.0
CE + SimCLR [35] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0 52.9 87.3
CSI [36] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5 64.6 87.0
SSD+ [37] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2 64.5 87.4
KNN+ [11] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4 55.1 88.3
CIDER [19] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3 31.1 93.0
Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3 26.6 96.9
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and precede the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations of our approach and the outlook of further work in the section of conclusions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the sull set of assumptions and a complete proof in the main manuscript and supplemental.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide full information about the replication experiments in the main manuscript and make our code and data publicly available.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We made our code and data publicly available on github as anonymous. The anonymous link is https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We present the details of the experiments in the supplemental material.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report results of experiments with statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the information about computer resources in the supplemental material.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We made our code and data publicly as anonymous in the review stage.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impacts of our work in the section of conclusions.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We have no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original paper that produced the code package or dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We discuss the details of our new dataset.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
-Adapting Multi-modal Large Language Model to Concept Drift in the Long-tailed Open World
+Adapting Multi-modal Large Language Model to Concept Drift From Pre-training Onwards
 Abstract
-Real-world data often exhibit extreme imbalances and out-of-distribution (OOD) instances, which significantly biases the model training.
-While it has been extensively studied in vision and language domains separately, the impact of long-tailed open worlds on multi-modal large language models (MLLMs) has been largely overlooked.
-In this paper, we first demonstrate the susceptibility and vulnerability of vision-language models to significant biases caused by tail drift and out-of-distribution (OOD) drift during both the pre-training and fine-tuning stages.
-To eliminate the bias from different sources, we integrate the tailed drift adaptation and OOD drift detection into a unified framework by extending the concept drift theory to multi-modal.
-Specifically, a T-distribution-based drift adapter is proposed to effectively mitigate the bias induced by the long-tailed problem, which also facilitates the model in distinguishing OOD data through explicit distribution modelling.
-Extensive experiments show significant improvements in our model’s ability to adapt to tailed drift and OOD drift.
-Moreover, it enhances the efficiency and accuracy of image-text alignment in vision language model pre-training, particularly in the long-tail open world scenario.
-Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world scenario, to validate our findings.
+Multi-modal Large Language Models (MLLMs) frequently face challenges from concept drift when dealing with real-world streaming data, wherein distributions change unpredictably.
+This mainly includes gradual drift due to long-tailed data and sudden drift from Out-Of-Distribution (OOD) data, both of which have increasingly drawn the attention of the research community.
+While these issues have been extensively studied in the individual domain of vision or language, their impacts on MLLMs in concept drift settings remain largely underexplored.
+In this paper, we reveal the susceptibility and vulnerability of Vision-Language (VL) models to significant biases arising from gradual drift and sudden drift, particularly in the pre-training.
+To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to the distribution unpredictable changes.
+Additionally, a T-distribution based drift adapter is proposed to effectively mitigate the bias induced by the gradual drift, which also facilitates the model in distinguishing sudden distribution changes through explicit distribution modeling.
+Extensive experiments demonstrate our method enhances the efficiency and accuracy of image-text alignment in the pre-training of VL models, particularly in the concept drift scenario.
+Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to long-tailed open world.
+Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world settings, to validate our findings.
 To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
 1 Introduction
-The explosive growth of data has precipitated a crisis for multi-modal large language models (MLLMs), which struggle to adapt to the long-tailed open world, encompassing the tailed drift and the out-of-distribution (OOD) drift.
-The tailed drift is brought by extremely imbalanced data, where the distribution of tailed categories shifts due to their inherent sparsity and noise, as well as the dominant head categories’ squeezing effect in the feature space.
-Meanwhile, OOD drift often occurs alongside the tailed drift and disrupts the underlying distribution of the model, as we continually encounter new concepts while navigating in an open world.
-Although the critical challenge of the long-tailed open world has been increasingly drawing attention in the realms of visual models [1] and language models [2], its impact on MLLMs, particularly vision-language (VL) models, has received scant attention.
-Therefore, we provide a comprehensive analysis of the impact of tail drift and OOD drift on VL models during both the pre-training and fine-tuning phases.
+The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift [1].
+Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolves because of their intrinsic sparsity and noise.
+Concurrently, sudden drift is mainly represented by OOD drift, as the model encounters new, previously unseen concepts, resulting in distributional shifts that disrupt its ability to generalize in an open-world context.
+While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models [2] and language models [3], their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored.
+In this work, we aim to bridge this gap by providing a systematic analysis of how tailed drift and OOD drift affect VL models during both pre-training and fine-tuning phases.
+Our findings highlight critical vulnerabilities of current VL models in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in dynamic, open-world environments.
 Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet [4] and the imbalanced dataset ImageNet-LT [2] is conducted.
@@
 It aligns image-text features for contrastive learning in pre-training.
-The desirable light-tailed property of the proposed T-distributed spherical metric (tSP) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift.
+The desirable light-tailed property of the proposed T-distributed spherical metric (Thp) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift.
 Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution.
-The proposed tSP distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings.
-And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the tSP distribution.
+The proposed T-hp distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings.
+And a simple non-parametric KNN is adopted to distinguish the OOD sample based on the T-hp distribution.
 Finally, we construct a group of multi-modal long-tailed open datasets to support our claims.
@@
   1. 1.
-    We are the pioneers in revealing the unexplored impacts of long-tailed open world to multi-modal large language models, especially in the image-text alignment in the pre-training and feature space allocation in the fine-tuning.
+    We are the pioneers in revealing the unexplored impacts of concept drift to multi-modal large language models, especially in the image-text alignment in the pre-training and feature space allocation in the fine-tuning.
 This allows future research to more comprehensively study the impact of defect data on MLLMs.
@@
 Assume that there are $N$ data streams corresponding to $N$ modalities, given a set of examples denoted as $S_{0,t}=\{S_{0},...,S_{i},...,S_{t}\}$, where $S_{i}=(s_{1},...,s_{j},...,s_{N})$ and $s_{j}=(X_{ij},y_{i})$ is one data instance from a single $j$-th data stream, $X_{ij}$ is the feature vector, $y_{i}$ is the label and $t$ is the timestamp of the data stream.
-$S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $F_{0,t}(S_{i})\neq F_{t+1,\infty}(S_{i})$, denoted as:
+$S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $P_{0,t}(S_{i})\neq P_{t+1,\infty}(S_{i})$, denoted as:
 $\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i})$ (1)
@@
 Furthermore, data drift predominantly stems from internal factors like data collection and processing, whereas concept drift typically arises from external factors, reflecting real-world changes.
-2.2 T-distributed Adapter for Long-Tailed Open World
-Within the concept drift framework, data distribution-based drift detection and adaptation have emerged as the dominant approaches [5, 6].
-To effectively mitigate the concept drift in the long-tail open world, we integrate directional statistics into distribution-based drift detection and adaptation, and propose a T-distributed adapter to alleviate it in the vision language model.
-2.2.1 Directional statistics
-Directional statistics primarily focus on the distribution of eigenvector angles, while neglecting the impact of eigenvector module lengths.
-Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set.
-A key idea in directional distribution is the tangent-normal decomposition.
-Any unit vector $x$ can be decomposed as:
-$x=t\mu+(1-t^{2})^{\frac{1}{2}}v,t\in[-1,1]$ (2)
-with $v\in\mathbb{S}^{d-2}$ a tangent to $\mathbb{S}^{d-1}$ at $\mu$ [7, 8], where $v$ and $t$ are independent and $v$ is uniform on $\mathbb{S}^{d-2}$.
-Thus, the intersection of $\mathbb{S}^{d-1}$ with the hyperplane through $t\mu$ and normal to $\mu$ is a $(d-2)$-dimensional sphere of radius $\sqrt{1-t^{2}}$, that $t$ has density as following:
-$p_{T}(t;d)\propto(1-t^{2})^{\frac{d-3}{2}},t\in[-1,1]$ (3)
-Therefore, through the marginal density $p_{T}$ and $p_{v}$, we can estimate the density of the entire spherical distribution.
-One prominent instance is the von Mises-Fisher distribution (vMF) [9], which can be interpreted as a probability distribution over the cosine similarity between a unit vector $x$ and a fixed mean direction $\mu$, following the density:
-$p_{X}(x;\mu,\kappa)\propto\exp{(\kappa\mu^{T}x)}$ (4)
-where $\kappa\geqslant 0$ denotes the concentration and $\exp$ represents the exponential function. Therefore, combined with the Eq. 2 and Eq. 3, the density of vMF is:
-$p(x)=C_{X}(\kappa,d)^{-1}\exp{(\kappa\mu^{T}x)},\quad x\sim\text{vMF}(\mu,\kappa)$ (5)
-$C_{X}(\kappa,d)=\frac{(2\pi)^{d/2}I_{d/2-1}(\kappa)}{\kappa^{d/2-1}}$
-where $I_{m}$ denotes the modified Bessel function of the first kind at order $m$.
-2.2.2 T-distributed Adapter
-Although the spherical nature of the vMF mitigates the bias induced by module length in long-tail open worlds, its exponential property, as evident in Eq.
-4, leads to a winner-take-all behaviour, where only a marginally higher logit than its competitors will dominate the eigenspace, receiving an exponentially disproportionate share of the eigenspace.
-As a consequence, the head categories, which possess abundant samples, exert a dominant influence over the entire model and squeeze the eigenspace of tailed categories, causing the crowding problem.
-Therefore, inspired by the T-SNE [10], we design a T-distribution based spherical metric (tSP), which follows the density:
-$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)}$ (6)
-where the unit feature vector $x\in\mathbb{S}^{d-1}$ and the center $\mu\in\mathbb{S}^{d-1}$. Through Eq. 3, we can get the marginal normalizer:
-$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (7)
-where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. 2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
+Figure 2: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks.
+Within the data streaming, a drift adaptation window slides to detect changes in data distribution and subsequently update the model, in both pre-training and fine-tuning.
+In the pre-training, the T-distributed adapter aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin.
+Coupled with the language model loss, they drive the training of all modules.
+In the downstream task, the image encoder and the text decoder are frozen out of training, with a linear projector fusing image-text features.
+Additionally, a mixture of experts module is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
+2.2 T-distributed Adapter for Concept Drift
+To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as
+$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)})$ (2)
+where $f^{(t)}$ denotes the vision language model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$.
+And the model is driven by the target metric $L$ continuously to adapt the drift in a given time period $[t,t+\tau]$.
+Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window [5, 6].
+Building upon this thinking, we integrate directional statistics into distribution-based drift detection and adaptation, proposing a T-distributed adapter to alleviate it in the vision language model.
+Firstly, we provide an overview of directional statistics in the Appendix LABEL:appendix:ds.
+Then, we will introduce the T-distributed adapter.
+Inspired by the T-SNE [7], we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
+$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})}$ (3)
+where $x^{(i)}\in\mathbb{S}^{d-1}$ denotes the unit feature vector, $\mu\in\mathbb{S}^{d-1}$ represents the center of category and $\kappa\geq 0$ symbolizes the concentration of the distribution, with higher values indicating a greater concentration around the center $\mu$.
+Accordingly, we can get the marginal normalizer:
+$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (4)
+where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. LABEL:eq:2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
 $N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$ (5)
-Thus, the probability density function of the proposed tSP distribution is as follows:
-$p(x)=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)},\quad x\sim\text{tSP}(\mu)$ (9)
-The detailed derivation process is provided in Appendix A.2.
+Thus, the probability density function of the proposed Thp distribution is as follows:
+$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu)$ (6)
+The detailed derivation process is provided in Appendix LABEL:appendix:Tsp.
 Figure 3: The proposed T-distributed spherical metric with various $\kappa$ and the classical vMF metric when $\kappa=1$.
-In terms of adapting to tailed concept drift, the tSP metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 2.
-The high kurtosis of tSP is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers.
-Formally, $L_{\text{tsp}}(\mu,x)=\frac{2}{\kappa(1-\mu^{T}x)+\epsilon}$ denotes the T-distributed spherical metric where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric.
+In terms of adapting to tailed concept drift, the Thp metric with a large concentration exhibits a light-tailed property, wherein the probability density function exhibits a faster rate of decay as the values increase, relative to the vMF metric, as illustrated in Figure 3.
+The high kurtosis of Thp is characterized that it yields high confidence only when the feature vector is sufficiently close to the center of the category, thereby minimizing the influence of head category samples on the tail category centers.
+Formally, $L_{\text{thp}}(\mu,x^{(i)})=\frac{2}{\kappa(1-\mu^{T}x^{(i)})+\epsilon}$ denotes the T-distributed metric on hypersphere, where $\epsilon$ is a non-zero value setting to $1$ to avoid the denominator of 0, and $L_{\text{vmf}}(\mu,x)=\exp(\kappa\mu^{T}{x})$ represents the vMF metric.
 Given an unit feature vector $x_{\text{head}}$ from the head categories, the gradient of the metric $L$ over the tailed category center $\mu_{\text{tail}}$ is $\frac{\partial L(\mu_{\text{tail}},x_{\text{head}})}{\partial\mu_{\text{tail}}}$.
-Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{tsp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$.
-Consequently, the light-tailed tSP distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
-Likewise, the tSP metric is directly applied to detect the OOD concept drift.
+Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{thp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$.
+Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
+Likewise, the Thp metric is directly applied to detect the OOD concept drift.
 A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace.
@@
 OOD), which does not impose any distributional assumption on the feature space.
-Here the distance is the tSP metric with respect to the k-th nearest neighbor.
-Figure 3: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks.
-By framing multi-modal heterogeneous data as concept drift data streams, we unify both tail drift and OOD drift.
-In the pre-training, the tSP metric aligns visual and textual feature space by image-text contrastive learning, with a large inter-class margin.
-Coupled with the language model loss, they drive the training of all the module.
-In the downstream task, the image encoder and the text decoder are frozen out of training, and two linear layers as the projector embed features into the hyperspherical eigenspace.
-The T-distributed adapter explicitly incorporates light-tail property into the feature space modelling, allowing it to effectively adapt tail drift, and performs OOD drift detection based on the distribution.
-2.3 T-distributed Vision Language Model for the Long-Tailed Open World
+Here the distance is the Thp metric with respect to the k-th nearest neighbor.
+2.3 T-distributed Vision Language Model for the Concept Drift
 As illustrated in Figure 2, our proposed vision language model contains two stages, the pre-training and the fine-tuning on the downstream task.
 Specifically, the classification task is chosen to visualize the impact of the bias caused by the long-tailed open world on the model.
-The multi-modal concept drift theory offers a unified framework for integrating tail drift adaptation and OOD drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
+The multi-modal concept drift theory offers a unified framework for integrating gradual drift adaptation and sudden drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
 The architecture of the proposed vision language model contains the image encoder, text encoder and text decoder.
@@
 It couples with the language modelling loss to drives the training of all three modules.
-In the downstream classification task, we leverage the T-distributed adapter to mitigate the long-tailed distribution of categories, thereby improving classification performance.
-Furthermore, by explicitly modelling the feature space, the T-distributed adapter enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples.
+In the downstream classification task, we leverage the T-distributed adapter as the router to distribute features to various FFNs as experts.
+Furthermore, by explicitly modelling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples.
 Besides, following the Blip [12], the image encoder and the text decoder are frozen out of training during fine-tuning.
 A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
-2.4 Building Multi-modal Datasets Group OpenMMlo for the Long-Tailed Open World
+2.4 Building Multi-modal Dataset OpenMMlo for the Long-Tailed Open World
 Due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to follow the long-tailed distribution in the open world.
@@
 Please describe in detail its size, location, color, and its relationship to the surroundings.".
-For more details about OpenMMlo, please refer to Appendix A.4.
+For more details about OpenMMlo, please refer to Appendix LABEL:appendix:details.
 3 Experiments
@@
 The model is trained on CIFAR100-LT [17] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [18], Places365 [19], LSUN [20], iSUN [21] and Texture [22].
-More detailed experimental implementations are given in Appendix A.3.
+More detailed experimental implementations are given in Appendix LABEL:appendix:expdetails.
 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
-We compare our proposed vision-language model with other specialized models to explicitly demonstrate its superior performance in long-tailed open-world scenarios.
-As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across all three large-scale datasets.
-The results validate the robustness of our vision language model against biases arising from tail drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
-Table 1: Evaluation results of long-tailed classification on ImageNet-LT, iNatualist2018 and Places-LT datasets with Vit-Large.
+We compare our proposed vision-language model with other models to explicitly demonstrate its superior performance in long-tailed open-world scenarios.
+As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across two large-scale datasets, namely the ImageNet-LT and iNaturalist 2018.
+Especially, training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip [23] pre-trained by the large WIT dataset.
+The results validate the robustness of our vision language model against biases arising from tailed drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
+Table 1: Evaluation results of long-tailed classification on ImageNet-LT and iNatualist2018.
 The best-performing models are highlighted in red.
@@
 Top-1 accuracy is applied to evaluate the performance of different methods.
-Methods ImageNet-LT iNaturalist 2018 Places-LT
-Many Medi. Few All Many Medi. Few All Many Medi. Few All
-MiSLAS [26] 62.0 49.1 32.8 51.4 - - - 70.7 - - - 38.3
-Bl-Softmax [27] 64.1 48.2 33.4 52.3 - - - 70.6 - - - 39.4
-LADE [28] 64.4 47.7 34.3 52.3 - - - 69.3 - - - 39.2
-ACE [29] - - - 56.6 - - - 72.9 - - - -
-RIDE [30] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6 - - - -
-PaCo [31] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2 36.1 47.9 35.3 41.2
-NCL [32] - - - 60.5 - - - 74.9 - - - 41.8
-MEM [33] 76.0 62.0 43.7 64.9 83.6 83.0 81.3 82.4 49.1 47.4 37.4 46.0
-Ours 80.3 71.1 37.2 65.6 84.3 83.1 80.7 82.6 61.6 53.3 32.6 46.9
-Furthermore, our model exhibits superior performance on the Many and Medium splits, while it slightly lags behind other models in the Few split.
-On one hand, our findings align with the scaling law observed in large models, where the performance exhibits a power law relationship with the size of the data.
-As the amount of data increases, the model’s feature extraction capability becomes more robust leading to improved performance.
-On the other hand, our goal is to alleviate the bias induced by tail drift, instead of averaging the feature space of each category and sacrificing head categories to enhance the performance of tail categories.
+ImageNet-LT iNaturalist 2018
+Methods Backbones Many Medium Few All Many Medium Few All
+Training from scratch
+cRT [24] 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
+LWS [24] 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
+MiSLAS [25] 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
+BALMS [26] 64.1 48.2 33.4 52.3 - - - 70.6
+LADE [27] 64.4 47.7 34.3 52.3 - - - 69.3
+ACE [28] - - - 56.6 - - - 72.9
+RIDE [29] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
+PaCo [30] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
+NCL [31] ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
+ViT [9] 73.7 46.5 15.6 52.4 65.4 55.3 50.9 54.6
+MAE [32] 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
+DeiT [33] 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
+LiVT [34] ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
+Ours 77.2 68.6 51.3 69.6 83.3 82.1 80.5 81.5
+Only Fine-tuning
+LPT [35] - - - - 62.1 76.2 79.3 76.1
+BALLAD [36] 79.1 74.5 69.8 75.7 - - - -
+Decoder [37] - - - 73.2 - - - 59.2
+VL-LTR [38] 84.5 74.6 59.3 77.2 - - - 81.0
+LIFT [39] ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
+Ours 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
+Compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks.
+Furthermore, to compare the current long-tailed methods, we the same setup as LIFT[39], i.e., using the pre-trained model of clip and only fine-tuning.
+The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data.
+Besides, although we are slightly behind LIFT[39] in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
 […]
-Methods SVHN LSUN iSUN Texture Average
-FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC
-ProxyAnchor [34] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0 65.0 86.0
-CE + SimCLR [35] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0 52.9 87.3
-CSI [36] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5 64.6 87.0
-SSD+ [37] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2 64.5 87.4
-KNN+ [11] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4 55.1 88.3
-CIDER [19] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3 31.1 93.0
-Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3 26.6 96.9
+Methods SVHN LSUN iSUN Texture
+FPR AUROC FPR AUROC FPR AUROC FPR AUROC
+ProxyAnchor [40] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
+CE + SimCLR [41] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
+CSI [42] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
+SSD+ [43] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
+KNN+ [8] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
+CIDER [16] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
+Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3
 In terms of OOD drift detection, our proposed vision language model, trained on CIFAR100-LT as an in-distribution dataset, demonstrates exceptional performance across four diverse OOD datasets, as shown in Table 2.
@@
 The proposed methodology offers the convenience of training and inference for large models.
+3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
 Table 3: Evaluation results of image-text alignment of different contrastive learning strategies in the stage of pre-training, from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories.
 The cosine metric is utilized to measure these distances, which is expressed as average degrees with standard deviation in brackets.
-We compare our proposed tSP with classical cosine loss, under balanced scenario (BL, ImageNet) and imbalanced scenarios (LT, ImageNet-LT), respectively.
+We compare our proposed Thp with classical cosine loss, under balanced scenario (BL, ImageNet) and imbalanced scenarios (LT, ImageNet-LT), respectively.
 Pre-training
@@
 LT / Cosine 49.2 ($\pm$4.95) 76.5 ($\pm$2.16) 80.7 ($\pm$1.90)
-LT / tSP 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
-3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
+LT / Thp 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
 Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training.
-As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/tSP.
+As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/Thp.
 It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios.
@@
 3.3.1 T-distributed Spherical Embedding in the Pre-training and Fine-tuning
-Table 4: Ablation evaluation results with or without the T-distributed adapter in the pre-training or fine-tuning.
-The ✓denotes the stage is trained with the T-distributed adapter.
-The results are based on the ImageNet-LT with the Vit-base.
-Top-1 accuracy (Acc) is used as the metric.
-Pre-training Fine-tuning Acc
-- - 53.5
-- 55.7
-- 60.0
-63.8
 Firstly, we conduct ablation experiments to verify the improved performance of the T-distributed adapter in the stage of pre-training and fine-tuning, respectively.
@@
 As a result, the T-distribution adapter’s impact is more pronounced in the fine-tuning stage compared to pre-training.
+Table 4: Ablation evaluation results with or without the T-distributed adapter in the pre-training or fine-tuning.
+The ✓denotes the stage is trained with the T-distributed adapter.
+The results are based on the ImageNet-LT with the Vit-base.
+Top-1 accuracy (Acc) is used as the metric.
+Pre-training Fine-tuning Acc
+- - 56.0
+- 58.7
+- 65.1
+69.6
 3.3.2 Various Concentration Parameter $\kappa$
@@
 $\kappa$ 4 16 64 128 Training
-Acc 61.3 63.8 59.9 57.2 62.3
+Acc 67.2 69.6 64.9 61.1 68.3
 In Table 5, setting the concentration parameter to $\kappa=16$ yields superior results.
 We argue that, a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model.
-In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the tSP metric.
+In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the Thp metric.
 In the context of concentration as a trainable parameter with the initialization of 16, there is a slight reduction in model performance, accompanied by a marginal increase of the concentration parameter to 16.37.
@@
 In future research, we will leverage concept drift theory to analyze the temporal, intensity, and regional aspects of the various drifts.
-Appendix A Appendix / supplemental material
-A.1 Related Works
-A.1.1 Multi-modal Large Language Model
-Large Language Models (LLMs) have recently significantly impacted the field of natural language processing.
-Through alignment techniques such as supervised learning and reinforcement learning with human feedback, LLMs can effectively generalize to perform a wide range of tasks, even with limited training data.
-A remarkable application of LLM is ChatGPT, which presents an amazing ability to interact with humans.
-OpenAI’s ChatGPT and GPT4 are prime examples of the impact that AI can have, and there have been extensive open-source efforts to replicate their success, such as OPT [38], BLOOM [39], PALM [40], LLaMA [41].
-Multi-modal large language models have further promoted the development of the vision-language model [42, 43, 44, 45, 46, 47, 48, 49].
-CLIP [42] was introduced to separately extract features from different encoders and combine them using contrastive learning.
-Building on CLIP, GLIP [43] was developed to learn object-level, language-aware, and semantic-rich visual representations, unifying object detection and phrase grounding for pre-training.
-Different from the contrastive method, Flamingo [44] aligned a pre-trained vision encoder and language model using gated cross-attention, demonstrated impressive few-shot learning capabilities.
-BLIP2 [45] was subsequently introduced, and it employed a Flan-T5 [50] along with a Q-Former to effectively align visual features with the language model.
-MiniGPT4 [46], the most recent development in the field is the PaLM-E model, which features 562 billion parameters and is designed to integrate real-world continuous sensor modalities into an LLM, thereby establishing a connection between real-world perceptions and human languages.
-Based on Visual Fundamental Models like BLIP mentioned above, Visual ChatGPT adopts ChatGPT as the central component for interacting with users.
-It integrates multiple visual foundation models and utilizes prompt engineering, also known as Prompt Manager, to instruct ChatGPT about the usage, input-output format, and capabilities of each foundation model.
-This enables ChatGPT to determine how to invoke these models to fulfill the user’s requirements.
-Besides, GPT-4V(ision)[51] and GPT-4O(mni) have recently shown unprecedented ability in understanding and processing an arbitrary mix of input images and texts.
-A.1.2 Long-tailed Open World
-In vision tasks, significant efforts have been devoted to mitigating the challenges posed by the long-tailed open world.
-Two prominent research directions have emerged: long-tailed classification under open-world settings, exemplified by approaches like OLTR++ [16, 1], LUNA [52], DALC [53], Open-sampling [54] and TLC [55], and OOD detection in long-tailed recognition, as seen in methods such as PASCL [56], EAT [57].
-OLTR++ [16, 1] proposed an ensemble algorithm, consisting of dynamic meta-embedding to improve the recognition of tail categories and active learning for open categories detection.
-LUNA [52] presented a distribution-sensitive loss to weigh more on the tail classes and a local-density-based metric to measure the novelty of OOD samples.
-DALC [53] designed an active distribution optimization algorithm for clustering, querying and classification to balance the classification bias.
-Open-sampling [54] rebalances class priors by sampling labels from a complementary distribution for each open-set instance, mitigating class imbalance.
-TLC [55] utilizes Dempster-Shafer Evidence Theory in a multi-expert framework for uncertainty estimation of tail and OOD samples.
-PASCL [56] applied supervised contrastive learning to explicitly boost the model to distinguish between tail-class in-distribution samples and OOD samples.
-EAT [57] introduces abstention classes for clear decision boundaries and augmenting tail classes with context-rich OOD data to focus on discriminative features.
-MCM [58] pioneers the integration of vision language models into OOD detection, enabling zero-shot OOD by aligning visual features with text concepts through a proposed maximum concept matching approach.
-Additionally, in the domain of the language model, [2] corroborates that large language models (LLMs) also struggle to learn long-tailed knowledge.
-While larger models are better at absorbing the long-tailed knowledge, they estimate that current models must be scaled by many orders of magnitude to reach competitive performance.
-Besides, [59] alleviates the long-tail problem in neural machine translation by quantifying token classification and sequence generation, and introduces an anti-focus loss that incorporates beam search inductive biases to better adapt model training to conditional text generation.
-A.1.3 Concept Drift
-In the review [4], the algorithms related to concept drift are categorized into three groups: error rate-based, data distribution-based and multiple hypothesis-based.
-Our proposed algorithm belongs to the distribution-based concept drift detection and adaptation method.
-Distribution-based concept drift algorithms not only accurately detect drift through explicit distributions but also analyze the drift to identify its happening timing, location, and severity.
-Besides, RBM-IM [60] proposes a novel trainable concept drift detector based on Restricted Boltzmann Machine, to slove the concept drift in multi-class imbalanced data streams.
-CALMID [61] propose a comprehensive active learning method for multiclass imbalanced streaming data with concept drift, including an ensemble classifier, a drift detector, and a variable threshold uncertainty strategy.
-DES-ICD [62] is a dynamic ensemble selection method for imbalanced data streams with concept drift.
-It considers local performances of base classifiers and addresses class imbalance using a novel synthetic minority oversampling technique.
-GOOD [63] develops an graph OOD benchmark, which explicitly distinguishes between covariate and concept shifts and design data splits that accurately capture these different shifts.
-A.1.4 Hyperspherical Distribution Modelling
-The Bayesian estimation of vMF mixture model with variational inference is addressed in [64].
-The learning task in VI consists of the optimization of the variational posterior distribution.
-Besides, a deep metric learning model for image classification and retrieval is presented in [65] , which utilizes the vMF distribution to define the loss function and introduces an effective alternative learning algorithm by updating class centers.
-The model captures global information in the embedding space and approximates the class distribution during training, leading to improved performance in image tasks.
-[66] extends the vMF distribution to regularize the intra-class feature distribution for imbalanced, small-scale and noisy data.
-[67] focus on using hyperspherical embedding to alleviate the crowding problem arisen by the imbalanced data.
-[19] utilizes hyperspherical embeddings for OOD detection in representation learning, consisting of two losses, a dispersion loss to increase angular distances between different class prototypes, and a compactness loss to ensure samples are closer to their respective class prototypes.
-A.2 The T-distributed Spherical Distribution
-Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set.
-The proposed T-distribution based spherical metric (tSP) follows the density:
-$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)}$ (12)
-where $x\in\mathbb{S}^{d-1}$, direction $\mu in\mathbb{S}^{d-1}$ and concentration $\kappa\in\mathbb{R}_{\geq 0}$.
-Let $T$ bet a random variable that denotes the dot-product $t=\mu^{T}x$, then $T=2Z-1$, with $Z\sim\text{Beta}(\alpha,\beta)$, where $\alpha=\frac{d-1}{2}$ and $\frac{d-3}{2}$.
-Proof.
-Given Eq. 3, the marginal distribution of the dot-product $t$ is
-$t\propto\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}$ (13)
-So, its normalizer is:
-$N_{T}(\kappa,d)$ $=\int_{\mathbb{S}^{d-1}}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}\mathrm{d}t$ (14)
-$=\int_{-1}^{1}\frac{1}{\kappa(1-t)}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-3}{2}}\mathrm{d}t$
-$=\frac{1}{\kappa}\int_{-1}^{1}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}\mathrm{d}t$
-Given the useful integral function:
-$\int(1+x)^{a}(1-x)^{b}\mathrm{d}x=2^{a+b+1}B_{\frac{x+1}{2}}(a+1,b+1)+C$ (15)
-So, its normalizer is:
-$N_{T}(\kappa,d)$ $=\frac{1}{\kappa}2^{d-3}(B_{1}(\frac{d-1}{2},\frac{d-3}{2})-B_{0}(\frac{d-1}{2},\frac{d-3}{2}))$ (16)
-$=\frac{1}{\kappa}2^{d-3}B(\frac{d-1}{2},\frac{d-3}{2})$
-The Beta function:
-$B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$ (17)
-So, the normalizer is
-$N_{T}(\kappa,d)=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (18)
-where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. It follows that the probability density function of the marginal distribution of the dot product is
-$p_{T}(t;\kappa,d)$ $=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}$ (19)
-$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}$
-$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(2z)^{\frac{d-1}{2}-1}(2-2z)^{\frac{d-3}{2}-1}$
-$=\frac{2}{\kappa}B(\alpha,\beta)^{-1}z^{\alpha-1}(1-z)^{\beta-1}$
-where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. ∎
-Due to the surface area of the hyper-sphere $\mathbb{S}^{d-1}$ is:
-$A_{d-1}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}$ (20)
-The T-distributed sperical distribution is expressed via the tangent normal decomposition as a joint distribution between $T\sim p_{T}{t;\kappa,d}$ and $V\sim\mathcal{U}(\mathbb{S}^{d-2})$.
-Since $T\perp\!\!\!\perp V$, the tSP normalizer $N_{x}(p,k)$ is the product of the normalizer of $p_{T}(t;\kappa,d)$ and the uniform distribution on $\mathbb{S}^{d-2}$ is:
-$N_{X}(\kappa,d)$ $=N_{T}(\kappa,d)\cdot A_{d-2}$ (21)
-$=2^{\alpha+\beta-1}B(\alpha,\beta)\frac{2\pi^{\beta}}{\kappa\Gamma(\beta)}$
-$=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$
-Thus,
-$p_{X}(x;\mu,\kappa)=N_{X}(\kappa,d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)}$ (22)
-A.3 Implementation Details
-For our language-guided image tokenizer, we leverage the strengths of both BERT [68] and ViT as our text encoder, text decoder and visual encoder, respectively.
-We employ ViT-Large as our visual encoder, which consists of 24 transformer encoder layers and an FFN intermediate size of 4,096.
-The input image size is set to $384\times 384$, with a patch size of $16\times 16$.
-The hidden dimensions of the ViT-Large are 1,024, with 16 attention heads.
-And, the number of parameters is about 307 million.
-Besides, we also use Vit-Base and ResNeXt-50 to perform ablation experiments.
-Vit-Base consists of 12 transformer encoder layers and an FFN intermediate size of 3,072.
-The hidden dimensions of the ViT-Base are 768, with 12 attention heads.
-The number of parameters is about 86 million.
-The input image size is set to $224\times 224$.
-In addition, ResNeXt-50 has 16 residual blocks with 50 layers.
-Each block has 3 convolutional layers with the kernel size of $3\times 3$, the stride of 1 and the padding of 1.
-The batch normalization and max pooling are utilized to connect the convolutional layers.
-The classification head hidden dimensions are 2,048.
-Additionally, BERT as the language model in our vision-language model, has 12 transformer layers with 768 hidden dimensions and 3,078 intermediate dimension.
-The number of attention heads is 12, with the input sequence length of 512.
-It has approximately 110 million parameters.
-(a) Sample in Training Set
-(b) Sample in Test Set
-(c) Sample in Open Set
-Figure 4: Samples of OpenMMlo in training set, test set and open set.
-In terms of the pre-training progress, the hyperparameters are presented in Table 6.
-We utilize the AdamW optimizer, which is configured with a cosine annealing schedule as the learning policy.
-The initial learning rate is set to $2\times 10^{-5}$, and the AdamW optimizer is employed with hyperparameters $\beta=(0.9,0.98)$.
-Additionally, we set the weight decay to 0.05 and the dropout rate to 0.1.
-During the first 1,000 warm-up steps, the learning rate increases to $2\times 10^{-5}$, and subsequently decays to $10^{-7}$.
-Unless otherwise specified, the pre-training of our vision language model consists of 800,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs, which takes approximately 7 days to complete.
-Table 6: The training hyperparameters of our vision language model.
-Pre-training
-Training Steps 400,000
-Warmup Steps 1,000
-Optimizer AdamW
-Learning Rate 1e-4
-Learning Rate Decay Cosine
-Adam $\beta$ (0.9, 0.98)
-Weight Decay 0.05
-Batch Size 50
-Fine-tuning
-Training Steps 18,000
-Warmup Steps 0
-Optimizer AdamW
-Learning Rate 2e-5
-Learning Rate Decay Cosine
-Adam $\beta$ (0.9, 0.98)
-Weight Decay 0.05
-Batch Size 400
-While in the fine-tuning on downstream task of classification, the initial learning rate is reduced to $10^{-6}$ without the warmup.
-The visual encoder and text decoder are frozen out of the training.
-Thus, the batch size can be increased to 400.
-The fine-tuning consists of 18,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs, which takes approximately 15 hours to complete.
-Other training parameters are as the same as the pre-training.
-When evaluating the performance of our VL model under the long-tailed open world, we use the top-1 accuracy metric on the downstream classification task.
-In particular, the categories are split into three groups: many-shot (with more than 100 training samples), medium-shot (with 20-100 training samples), and few-shot (with fewer than 20 training samples).
-The Top-1 accuracies are computed for each group to evaluate the performance of mitigating the bias introduced by the long-tail distribution, respectively.
-Furthermore, in order to assess the capability of detecting the OOD drift, we employ two metrics: FPR95 which measures the false positive rate of OOD samples when the true positive rate of ID samples reaches 95%, and AUROC providing the area under the receiver operating characteristic curve.
-Besides, cosine distance is exploited to measure the distances between features and centers in the feature space of the VL model.
-A.4 Building Multi-modal Long-Tailed OOD Datasets Group OpenMMlo
-Figure 4 showcases the samples utilized for training and validation in our study.
-To intuitively verify the impact of long-tail open world scenarios on multi-modal large language models, we employ classification as our downstream task.
-When matching images and texts, we strategically mask words that are directly related to category names.
-This approach ensures the accuracy and reliability of our experimental results.
-As depicted in Figure 4, comprehensive descriptions of the image are provided through long-form text, encompassing details such as size, position, color, relationships, and other relevant information about the objects present in the image.
-This ensures a detailed and information-rich depiction of the visual content.
-We have publicly released the datasets used for training and validation, as well as the original unmasked datasets.
-A.5 More Experiment Results
-A.5.1 Various Image Encoder
-Table 7: Ablation evaluation results of various image encoders, including the ResNeXt50, Vit-base and Vit-Large. The results are based on the ImageNet-LT. Top-1 accuracy is used as the metric.
-Methods ResNeXt-50 Vit-Base Vit-Large
-CE 45.2 53.5 59.7
-$\tau$-normalized [69] 49.0 57.1 62.7
-cRT [69] 49.3 58.1 64.0
-LWS [69] 48.2 57.9 63.3
-MEM [69] 49.6 58.9 64.9
-Ours 52.2 63.8 65.6
-Beyond leveraging Vit-Large as the backbone, we also explore alternative image encoders in ablation experiments, namely Vit-Base, Vit-large and ResNeXt-50 to further dissect the model’s performance.
-We conducted ablation experiments on various image encoders to demonstrate the effectiveness of our method across different backbone architectures, as illustrated in Table 7.
-We select various backbones consisting of ResNeXt50, Vit-base and Vit-large, with different model size.
-The results confirm that our method consistently achieves superior performance across various backbone networks.
-A.5.2 More Results
-In this section, we provide more results of OOD detection.
-[…]
-Methods SVHN LSUN iSUN Texture Average
-FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC
-MSP [70] 78.9 79.8 83.5 75.3 84.6 74.5 86.5 72.5 84.4 74.2
-ODIN [71] 70.2 84.9 76.4 80.1 79.5 79.2 85.3 75.2 82.2 75.2
-Mahalanobis [72] 87.1 80.6 84.2 79.4 83.2 78.8 61.7 84.9 84.6 73.9
-Energy [73] 66.9 85.3 59.8 86.7 66.5 84.5 79.0 80.0 81.4 76.4
-GODIN [74] 74.6 84.0 93.3 67.2 94.3 65.3 86.5 69.4 89.1 69.0
-LogitNorm [75] 59.6 90.7 81.1 83.0 84.2 80.8 86.6 75.6 80.3 78.6
-ProxyAnchor [34] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0 65.0 86.0
-CE + SimCLR [35] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0 52.9 87.3
-CSI [36] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5 64.6 87.0
-SSD+ [37] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2 64.5 87.4
-KNN+ [11] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4 55.1 88.3
-CIDER [19] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3 31.1 93.0
-Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3 26.6 96.9
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and precede the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss limitations of our approach and the outlook of further work in the section of conclusions.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: We provide the sull set of assumptions and a complete proof in the main manuscript and supplemental.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We provide full information about the replication experiments in the main manuscript and make our code and data publicly available.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We made our code and data publicly available on github as anonymous. The anonymous link is https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We present the details of the experiments in the supplemental material.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We report results of experiments with statistical significance.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide the information about computer resources in the supplemental material.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We made our code and data publicly as anonymous in the review stage.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We discuss the broader impacts of our work in the section of conclusions.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We have no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We cite the original paper that produced the code package or dataset.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We discuss the details of our new dataset.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We do not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 In this paper, we reveal the susceptibility and vulnerability of Vision-Language (VL) models to significant biases arising from gradual drift and sudden drift, particularly in the pre-training.
-To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to the distribution unpredictable changes.
+To effectively address these challenges, we propose a unified framework that extends concept drift theory to the multi-modal domain, enhancing the adaptability of the VL model to unpredictable distribution changes.
 Additionally, a T-distribution based drift adapter is proposed to effectively mitigate the bias induced by the gradual drift, which also facilitates the model in distinguishing sudden distribution changes through explicit distribution modeling.
 Extensive experiments demonstrate our method enhances the efficiency and accuracy of image-text alignment in the pre-training of VL models, particularly in the concept drift scenario.
-Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to long-tailed open world.
-Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open world settings, to validate our findings.
-To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/Anonymous0Knight/ConceptDriftMLLMs.
+Moreover, various downstream tasks exhibit significant improvements in our model’s ability to adapt to the long-tailed open world.
+Furthermore, we create a set of multi-modal datasets called OpenMMlo, specifically tailored for the long-tailed open-world setting, to validate our findings.
+To foster the development of the multi-modal community, we have made both OpenMMlo datasets and our code publicly available at: https://github.com/XiaoyuYoung/ConceptDriftMLLMs.
 1 Introduction
-The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift [1].
-Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolves because of their intrinsic sparsity and noise.
+The rapid expansion of data availability has created significant challenges for multi-modal large language models (MLLMs), particularly in addressing concept drift, which predominantly manifests as gradual drift and sudden drift Lu et al.
+(2019).
+Among them, tailed drift represents a classic illustration of gradual drift, emerging due to severe data imbalance, where the distributions of long-tail categories evolve because of their intrinsic sparsity and noise.
 Concurrently, sudden drift is mainly represented by OOD drift, as the model encounters new, previously unseen concepts, resulting in distributional shifts that disrupt its ability to generalize in an open-world context.
-While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models [2] and language models [3], their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored.
+While the issues of long-tailed recognition and concept drift in open-world settings have been extensively studied in visual models Liu et al.
+(2022b) and language models Kandpal et al.
+(2023), their impact on MLLMs, particularly vision-language (VL) models, remains largely unexplored.
 In this work, we aim to bridge this gap by providing a systematic analysis of how tailed drift and OOD drift affect VL models during both pre-training and fine-tuning phases.
 Our findings highlight critical vulnerabilities of current VL models in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in dynamic, open-world environments.
-Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet [4] and the imbalanced dataset ImageNet-LT [2] is conducted.
+Pre-training: As illustrated in Figure 1(a), a comparison of the VL model trained on the balanced dataset ImageNet Russakovsky et al.
+(2015b) and the imbalanced dataset ImageNet-LT Liu et al.
+(2022b) is conducted.
 Due to the implicit feature centers of each category, we approximate them by averaging unit image and text features obtained by samples on the test set.
@@
 Figure 1: The impacts of tailed drift and OOD drift on the vision language model in the stages of pre-training and fine-tuning, respectively.
-(a) In terms of the pre-training, we visualize the alignment results on both a balanced dataset (denoted as BL) and an imbalanced dataset (as LT).
-The cosine metric is used to measure the distances between unit image and text features across various categories including OOD samples, which is expressed as degrees.
+(a) In terms of the pre-training, we visualize the alignment results pre-trained on both a balanced dataset (denoted as BL) and an imbalanced dataset (as LT) without OOD samples, under the same balanced test set.The cosine metric is used to measure the distances between unit image and text features across various categories including OOD samples, which is expressed as degrees.
 A smaller degree indicates a higher level of similarity between the features.
 Thus, it provides a feature-level visualization of the intra-class compactness and inter-class separability in the vision language model.
-(b) In the context of fine-tuning, the mutual cosine distance between the centers of each category in the classifier is directly visualized to illustrate the feature space of the classifier, denoted as blue bars.
+(b) In the context of fine-tuning in imbalance datasets, the mutual cosine distance between the centers of each category in the classifier is directly visualized to illustrate the feature space of the classifier, denoted as blue bars.
 Besides, the average cosine distance between each category center and OOD samples is calculated, which is represented as orange bars.
@@
 Moreover, the average cosine distance between each category and unit features extracted by OOD samples is applied to reveal the OOD separability as denoted as orange bars in Figure 1(b).
-Since head categories occupies most of the feature space, OOD samples are closer to the center of the head categories compared to the tail categories, implying that in the stage of fine-tuning, it is difficult for the VL model to distinguish between OOD samples under imbalanced scenarios.
-To effectively address tailed drift and OOD drift within a unified framework, which often occur simultaneously, we encapsulate them using the concept drift theory.
+Since head categories occupy most of the feature space, OOD samples are closer to the center of the head categories compared to the tail categories, implying that in the stage of fine-tuning, it is difficult for the VL model to distinguish between OOD samples under imbalanced scenarios.
+To effectively address tailed drift and OOD drift within a unified framework, which often occurs simultaneously, we encapsulate them using the concept drift theory.
 Therefore, summarizing the above challenges of vision language models in the long-tailed open world, it raises the important question:
@@
 The desirable light-tailed property of the proposed T-distributed spherical metric (Thp) prevents the compression of tailed categories and mitigates the crowding of feature space caused by tailed drift.
-Besides, in fine-tuning, the adapter projects features to decision boundary and detects OOD samples at feature-level based on the underlying distribution.
+Besides, in fine-tuning, the adapter projects the features to the decision boundary and detects OOD samples at the feature level based on the underlying distribution.
 The proposed T-hp distribution explicitly models the feature space with concrete feature centers, optimizing large inter-class margins and yielding more desirable hyperspherical embeddings.
@@
   4. 4.
-    We build a group of multi-modal datasets OpenMMlo under the long-tailed open world, by extending existing image-based long-tailed open datasets.
+    We build a group of multi-modal datasets OpenMMlo under the long-tailed open world by extending existing image-based long-tailed open datasets.
 It contains about 740k image-caption pairs with related category annotations.
@@
 2.1 Multi-modal Concept Drift Theory
-Concept drift is a phenomenon in which the statistical properties of a target domain change over time in an arbitrary way [1].
+Concept drift is a phenomenon in which the statistical properties of a target domain change over time in an arbitrary way Lu et al.
+(2019).
 Formally, given a set of examples denoted as the data stream $S_{0,t}=\{d_{0},...,d_{t}\}$, where $d_{i}=(X_{i},y_{i})$ is one data instance, $X_{i}$ and $y_{i}$ respectively denote the feature vector and the label, and $t$ represents the timestamp of the instance in the data stream.
@@
 In the context of multi-modal vision language models, we extend the concept drift theory from a single data stream to multiple data streams.
-Each data modal is associated with a distinct data stream.
+Each modality is associated with a distinct data stream.
 Thereby, the multi-modal concept drift framework can robustly handle the complex, heterogeneous data distributions inherent to vision language models.
@@
 $S_{0,t}$ follows a certain distribution $F_{0,t}(S_{i})$, the multi-modal concept drift occurs at timestamp $t+1$, if $P_{0,t}(S_{i})\neq P_{t+1,\infty}(S_{i})$, denoted as:
-$\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i})$ (1)
-Remark 2.1.
-Differences: Concept Drift vs.
-Data Drift (Covariate Drift) Data drift entails changes solely in the distribution of inputs $P(x)$, while concept drift involves alterations in both input and output distributions, i.e., $P(x)$ and $P(y)$, leading to changes in the decision boundary.
-Furthermore, data drift predominantly stems from internal factors like data collection and processing, whereas concept drift typically arises from external factors, reflecting real-world changes.
-Figure 2: The workflow of our methodology, consisting of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks.
+$\exists t:P_{t}(S_{i})\neq P_{t+1}(S_{i}).$ (1)
+Figure 2: The workflow of our methodology, which consists of two stages: the pre-training of the vision-language model and the fine-tuning on downstream tasks.
 Within the data streaming, a drift adaptation window slides to detect changes in data distribution and subsequently update the model, in both pre-training and fine-tuning.
@@
 In the downstream task, the image encoder and the text decoder are frozen out of training, with a linear projector fusing image-text features.
-Additionally, a mixture of experts module is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
+Additionally, a mixture of expert modules is leveraged with the T-distributed adapter as the router, allowing it to effectively adapt tail drift and perform OOD drift detection based on the distribution.
 2.2 T-distributed Adapter for Concept Drift
-To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as
-$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)})$ (2)
+To adapt the vision language model to concept drift, it is essential to adapt the model to align with the evolving data distribution, which can be formally defined as:
+$\min_{f^{(t)},f^{(t+1)},...,f^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(f^{(i)}(x^{(i)}),y^{(i)}),$ (2)
 where $f^{(t)}$ denotes the vision language model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$.
 And the model is driven by the target metric $L$ continuously to adapt the drift in a given time period $[t,t+\tau]$.
-Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window [5, 6].
+Thus, one prevalent method for detecting and adapting to concept drift involves designing metrics based on data distributions that can effectively counteract the impacts of sudden and gradual changes within the time window Jiao et al.
+(2022); Yu et al.
+(2024).
 Building upon this thinking, we integrate directional statistics into distribution-based drift detection and adaptation, proposing a T-distributed adapter to alleviate it in the vision language model.
-Firstly, we provide an overview of directional statistics in the Appendix LABEL:appendix:ds.
+Firstly, we provide an overview of directional statistics in the Appendix A.2.1.
 Then, we will introduce the T-distributed adapter.
-Inspired by the T-SNE [7], we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
-$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})}$ (3)
+Inspired by the T-SNE Van der Maaten & Hinton (2008), we design a T-distribution based metric in hypersphere (T-hp), which follows the density:
+$p_{X}(x^{(i)})\propto\frac{2}{\kappa(1-\mu^{T}x^{(i)})},$ (3)
 where $x^{(i)}\in\mathbb{S}^{d-1}$ denotes the unit feature vector, $\mu\in\mathbb{S}^{d-1}$ represents the center of category and $\kappa\geq 0$ symbolizes the concentration of the distribution, with higher values indicating a greater concentration around the center $\mu$.
 Accordingly, we can get the marginal normalizer:
-$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ (4)
-where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. LABEL:eq:2, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
-$N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}$ (5)
+$N_{T}(\kappa,d)=\int_{\mathbb{S}^{d-1}}{\frac{2}{1-\kappa\mu^{T}x^{(i)}}}\mathrm{d}x=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)},$ (4)
+where $\alpha=\frac{d-1}{2}$, $\beta=\frac{d-3}{2}$, and $\Gamma(\cdot)$ represents the gamma function. Combined with Eq. 9 in Appendix A.2.1, the normalizer $N_{X}(d)$ of density $p_{X}(x;\mu)$ is:
+$N_{X}(\kappa,d)=N_{T}(\kappa,d)\cdot A_{d-2}=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}.$ (5)
 Thus, the probability density function of the proposed Thp distribution is as follows:
-$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu)$ (6)
-The detailed derivation process is provided in Appendix LABEL:appendix:Tsp.
+$p(x^{(i)})=N_{X}(d)^{-1}\frac{2}{\kappa(1-\mu^{T}x^{(i)})},\quad x^{(i)}\sim\text{Thp}(\mu).$ (6)
+The detailed derivation process is provided in Appendix A.2.
 Figure 3: The proposed T-distributed spherical metric with various $\kappa$ and the classical vMF metric when $\kappa=1$.
@@
 Due to $\mu^{T}_{\text{tail}}x_{\text{head}}\in[-1,1]$, when $\kappa\geqslant 1$, it is readily obtain that $\frac{\partial L_{\text{thp}}}{\partial\mu_{\text{tail}}}<\frac{\partial L_{\text{vmf}}}{\partial\mu_{\text{tail}}}$.
-Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by tailed concept drift.
+Consequently, the light-tailed Thp distribution effectively counteracts the squeezing of tail categories caused by an overwhelming number of head samples, thereby alleviating the bias induced by the tailed concept drift.
 Likewise, the Thp metric is directly applied to detect the OOD concept drift.
 A sample with a unit feature vector $x$ is deemed out-of-distribution if it lies at a relatively large distance from the in-distribution (ID) data in the eigenspace.
-Following [8], a simple non-parametric KNN is adopted to partition the data into two sets (ID vs.
+Following Sun et al.
+(2022), a simple non-parametric KNN is adopted to partition the data into two sets (ID vs.
 OOD), which does not impose any distributional assumption on the feature space.
@@
 The multi-modal concept drift theory offers a unified framework for integrating gradual drift adaptation and sudden drift detection, where heterogeneous image and text inputs are treated as distinct data streams.
-The architecture of the proposed vision language model contains the image encoder, text encoder and text decoder.
-With an input image $I_{i}\in\mathbb{R}^{H\times W}$, the visual features $x_{\text{img}}$ are extracted by the image encoder $E_{\text{img}}$ and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
-$x_{\text{img}}=P_{\text{norm}}(E_{\text{img}}(I_{i}))\in\mathbb{R}^{n\times d}$ (7)
+The proposed vision language model follows the encoder-decoder mixture architecture of the Blip Li et al.
+(2022c), containing an image encoder, a text encoder and an image-grounded text decoder.
+With an input image $I_{i}\in\mathbb{R}^{H\times W}$, the visual features $x_{\text{img}}$ are extracted by the image encoder $E_{\text{img}}$ and further projected to the spherical eigenspace by the $L_{2}$ normalizer $P_{\text{norm}}$:
+$x_{\text{img}}=P_{\text{norm}}(E_{\text{img}}(I_{i}))\in\mathbb{R}^{n\times d},$ (7)
 where $n$ is the number of visual features and $d$ represents the feature dimension.
-The image encoder $E_{\text{img}}$ can be any common visual backbones, such as Vit-Base [9], Vit-Large [9] and ResNeXt-50 [10].
-In terms of the text encoder $E_{\text{txt}}$, with a processed input text sequence $T_{i}$, text features are extracted by the language encoder and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
-$x_{\text{txt}}=P_{\text{norm}}(E_{\text{txt}}(T_{i}))\in\mathbb{R}^{n\times d}$ (8)
-where $n$ is the number of input tokens and $d$ represents the feature dimension. In our case, Bert [11] is used as the language encoder.
-In the pre-training of the vision language model, the T-distributed adapter aligns image and text features by contrastive learning.
+The image encoder $E_{\text{img}}$ can be any common visual backbones, such as Vit-Base Dosovitskiy et al.
+(2021), Vit-Large Dosovitskiy et al.
+(2021) and ResNeXt-50 Xie et al.
+(2017).
+In terms of the text encoder $E_{\text{txt}}$, with a processed input text sequence $T_{i}$, text features are extracted by the language encoder and further projected to the spherical eigenspace by the $L_{2}$ normalizer $P_{\text{norm}}$:
+$x_{\text{txt}}=P_{\text{norm}}(E_{\text{txt}}(T_{i}))\in\mathbb{R}^{n\times d},$ (8)
+where $n$ is the number of input tokens and $d$ represents the feature dimension.
+In our case, Bert Devlin et al.
+(2019a) is used as the language encoder, where a [CLS] token is added to the start of the text input for sentence summarization.
+Additionally, an image-grounded text decoder is employed to produce a textual description corresponding to a provided image.
+Utilizing the input visual features $x_{img}$ and text features $x_{txt}$, we initially create fused multi-modal representations by merging the image and text feature embeddings.
+These combined features act as the keys and values within the cross-attention blocks in the image-grounded text decoder.
+Through conditioning on the already predicted partial sequence $y_{i<j}$, the decoder iteratively forecasts the token at position $j$, effectively producing textual descriptions corresponding across modalities.
+In the pre-training of the vision language model, the T-distributed adapter aligns image and text encoders by contrastive learning.
 It seeks to align visual and textual transformer feature spaces by promoting similar representations for positive pairs and dissimilar representations for negative pairs.
 More importantly, our approach circumvents model bias stemming from the long-tailed distribution of data.
-It couples with the language modelling loss to drives the training of all three modules.
+Specifically, given a mini-batch with $N$ image-text feature pairs, we calculate the $N\times N$ Thp similarity of the cross between image and text features.
+$N$ correct pairs are recognized as positive samples to maximize the Thp similarity, whereas the rest of $N^{2}-N$ are negative samples to minimize the similarity.
+And we follow the ALBEF Li et al.
+(2021) to use soft labels from a momentum encoder as training targets to account for the potential positives in the negative pairs.
+Additionally, coupled with the T-distributed adapter, language modeling loss is utilized to activate the image-grounding text encoder for generating coherent and detailed captions based on the image, further propelling the training of all three modules.
+Driven by language modeling loss, the model is trained to optimize a cross-entropy loss with label smoothing, to maximize the likelihood of the generated text in an autoregressive manner.
 In the downstream classification task, we leverage the T-distributed adapter as the router to distribute features to various FFNs as experts.
-Furthermore, by explicitly modelling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples.
-Besides, following the Blip [12], the image encoder and the text decoder are frozen out of training during fine-tuning.
+Furthermore, by explicitly modeling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples.
+Besides, following the Blip Li et al.
+(2022c), the image encoder and the text decoder are frozen out of training during fine-tuning.
 A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
 2.4 Building Multi-modal Dataset OpenMMlo for the Long-Tailed Open World
-Due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to follow the long-tailed distribution in the open world.
+As the parameters of large models continue to expand, the demand for extensive training data also escalates.
+However, due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to be balanced in an open world, while cleaning the data requires huge costs.
+Thus, our aspiration is for the model to adeptly acclimate to the imbalanced dataset by itself, acquiring abundant knowledge with more and more data but not exhibiting bias.
+In this context, a more realistic training dataset for vision language models is required to validate their potential to be trained under the long-tailed open world.
 Recognizing the demand for higher-quality multi-modal data with long-tailed distribution in an open world, we developed a group of datasets called Open Multi-modal Long-Tailed OOD Datasets (OpenMMlo).
-We extend the open-source datasets, namely ImageNet-LT [13], iNatualist2018 [14] and Places-LT [13].
+We extend the open-source datasets, namely ImageNet-LT Liu et al.
+(2019), iNatualist2018 Van Horn et al.
+(2018) and Places-LT Liu et al.
+(2019).
 ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category.
@@
 The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes.
-We use the InstructBLIP [15] to generate the related caption of the image, with the prompt of "What does this picture describe?
-Please describe in detail its size, location, color, and its relationship to the surroundings.".
-For more details about OpenMMlo, please refer to Appendix LABEL:appendix:details.
+We use the InstructBLIP Dai et al.
+(2023) to generate the related caption of the image, with the prompt of ”What does this picture describe?
+Please describe in detail its size, location, color, and its relationship to the surroundings.”.
+And, we define long-tailed data in image-caption pairs according to the image categories, which are provided in open-source image datasets.
+Concerning related captions based on images, we counted the word frequencies and found that their distribution is similar to the image categories distribution, which is imbalanced.
+For more details about OpenMMlo, please refer to Appendix A.4.
 3 Experiments
@@
 Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift.
-The constructed long-tailed multi-modal datasets OpenMMlo is utilized for training and validating.
-In terms of the OOD drift detection, we follow the setting of CIDER [16].
-The model is trained on CIFAR100-LT [17] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [18], Places365 [19], LSUN [20], iSUN [21] and Texture [22].
-More detailed experimental implementations are given in Appendix LABEL:appendix:expdetails.
+The constructed long-tailed multi-modal dataset OpenMMlo is utilized for training and validating.
+In terms of the OOD drift detection, we follow the setting of CIDER Ming et al.
+(2023).
+The model is trained on CIFAR100-LT Krizhevsky & Hinton (2009) with an imbalance ratio of 100, and validated on external OOD datasets including SVHN Netzer et al.
+(2011), Places365 Zhou et al.
+(2017), LSUN Yu et al.
+, iSUN Xu et al.
+and Texture Cimpoi et al.
+(2014).
+More detailed experimental implementations are given in Appendix A.3.
 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
@@
 As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across two large-scale datasets, namely the ImageNet-LT and iNaturalist 2018.
-Especially, training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip [23] pre-trained by the large WIT dataset.
+To ensure a more equitable comparison, we opted to conduct pre-training using ImageNet and iNaturalist datasets separately, rather than pre-training the entire OpenMMlo.
+Besides, it is worth noting that training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip Radford et al.
+(2021) pre-trained by the large WIT dataset.
 The results validate the robustness of our vision language model against biases arising from tailed drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
+As shown in Table 1, compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks.
+Furthermore, to compare the current long-tailed methods, we apply the same setup as LIFT Shi et al.
+(2024), i.e., using the pre-trained model of the clip and only fine-tuning.
+Only fine-tuning means that the method does not pre-train the model on the long-tail dataset, while directly using the parameters of CLIP pre-trained on high-quality and large-scale WIT dataset.
+And they only fine-tune the model on long-tailed datasets.
+It is worth noting that, most vision language models only focus on the impact of tailed drift in the fine-tuning, such as LPT DONG et al.
+(2023), BALLAD Ma et al.
+(2021), Decoder Wang et al.
+(2024), VL-LTR Tian et al.
+(2022) and LIFT Shi et al.
+(2024).
+We are the pioneers in revealing the unexplored impacts of concept drift from pre-training onwards.
+The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data.
+Besides, although we are slightly behind LIFT Shi et al.
+(2024) in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
 Table 1: Evaluation results of long-tailed classification on ImageNet-LT and iNatualist2018.
 The best-performing models are highlighted in red.
-Many, Medi.
-and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples).
+Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples).
 Top-1 accuracy is applied to evaluate the performance of different methods.
+Additionally, $\dagger$ means the model is trained with the resolution of $384\times 384$.
+Besides, ZS denotes the zero-shot results of the CLIP model, LP represents the linear probing results, and FT means the fine-tuning results.
 ImageNet-LT iNaturalist 2018
@@
 Training from scratch
-cRT [24] 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
-LWS [24] 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
-MiSLAS [25] 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
-BALMS [26] 64.1 48.2 33.4 52.3 - - - 70.6
-LADE [27] 64.4 47.7 34.3 52.3 - - - 69.3
-ACE [28] - - - 56.6 - - - 72.9
-RIDE [29] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
-PaCo [30] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
-NCL [31] ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
-ViT [9] 73.7 46.5 15.6 52.4 65.4 55.3 50.9 54.6
-MAE [32] 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
-DeiT [33] 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
-LiVT [34] ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
-Ours 77.2 68.6 51.3 69.6 83.3 82.1 80.5 81.5
+cRT Kang et al. (2019) 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
+LWS Kang et al. (2019) 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
+MiSLAS Zhong et al. (2021) 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
+BALMS Ren et al. (2020) 64.1 48.2 33.4 52.3 - - - 70.6
+LADE Huang et al. (2016) 64.4 47.7 34.3 52.3 - - - 69.3
+ACE Cai et al. (2021) - - - 56.6 - - - 72.9
+RIDE Wang et al. (2020) 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
+PaCo Cui et al. (2021) 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
+NCL Li et al. (2022b) ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
+ViT Dosovitskiy et al. (2021) 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
+MAE He et al. (2022) 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
+DeiT Touvron et al. (2022) 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
+LiVT Xu et al. (2023) 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
+LiVT $\dagger$ Xu et al. (2023) ViT-B/16 76.4 59.7 42.7 63.8 83.2 81.5 79.7 81.0
+Ours 76.4 66.2 48.9 67.6 82.5 79.8 77.1 78.9
+[…]
 Only Fine-tuning
-LPT [35] - - - - 62.1 76.2 79.3 76.1
-BALLAD [36] 79.1 74.5 69.8 75.7 - - - -
-Decoder [37] - - - 73.2 - - - 59.2
-VL-LTR [38] 84.5 74.6 59.3 77.2 - - - 81.0
-LIFT [39] ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
-Ours 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
-Compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks.
-Furthermore, to compare the current long-tailed methods, we the same setup as LIFT[39], i.e., using the pre-trained model of clip and only fine-tuning.
-The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data.
-Besides, although we are slightly behind LIFT[39] in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
+CLIP+ZS Radford et al. (2021) 82.0 70.4 69.6 70.5 9.9 5.3 4.6 5.5
+CLIP+LP Radford et al. (2021) 87.3 65.1 19.0 67.4 62.4 7.1 0.1 10.0
+CLIP+FT Radford et al. (2021) 83.0 65.0 39.9 68.5 79.4 67.6 59.1 65.4
+LPT DONG et al. (2023) - - - - 62.1 76.2 79.3 76.1
+BALLAD Ma et al. (2021) 79.1 74.5 69.8 75.7 - - - -
+Decoder Wang et al. (2024) - - - 73.2 - - - 59.2
+VL-LTR Tian et al. (2022) 84.5 74.6 59.3 77.2 - - - 81.0
+LIFT Shi et al. (2024) ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
+Ours 79.5 76.5 74.1 77.2 83.5 82.2 80.7 81.7
+Ours $\dagger$ 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
+Beyond that, we compare our method with the CLIP Radford et al.
+(2021) under zero-shot, linear probing and fine-tuning, where CLIP results are from the Decoder Wang et al.
+(2024).
+Based on the zero-shot results, it is evident that CLIP, even trained on large-scale and high-quality WIT datasets, struggles to address the issue of tailed drift.
+CLIP only achieves 5.5% accuracy on iNaturalist2018, and the accuracy variance between many-shot and medium-shot scenarios is 11.6% on ImageNet-LT.
+Our method significantly outperforms CLIP in dealing with tailed drift, especially on iNaturalist2018.
+It also indicates that training on a high-quality balanced dataset alone cannot effectively mitigate the bias induced by long-tail drift.
+Furthermore, the results of linear probing and fine-tuning demonstrate that imbalanced datasets can induce pronounced tailed drift in MLLMs, ultimately degrading model performance.
+The CLIP accuracy in Few-shot is only 39.9% under fine-tuning on ImageNet-LT, much lower than the 69.6% under zero-shot.
+It further verifies the challenges brought by imbalanced data in the training of MLLMs and the superiority of our method in adapting the MLLM to concept drift from pre-training onwards.
 […]
@@
 FPR AUROC FPR AUROC FPR AUROC FPR AUROC
-ProxyAnchor [40] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
-CE + SimCLR [41] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
-CSI [42] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
-SSD+ [43] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
-KNN+ [8] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
-CIDER [16] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
+ProxyAnchor Kim et al. (2020) 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
+CE + SimCLR Winkens et al. 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
+CSI Tack et al. (2020) 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
+SSD+ Sehwag et al. (2021) 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
+KNN+ Sun et al. (2022) 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
+CIDER Ming et al. (2023) 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
 Ours 8.3 98.7 20.3 97.5 32.5 95.2 45.1 96.3
@@
 The proposed methodology offers the convenience of training and inference for large models.
+Table 3: Evaluation results of generalization on ImageNet-Sketch Wang et al.
+(2019) with ImageNet Russakovsky et al.
+(2015a) as the source dataset.
+We compare our methods with other VL models, including zero-shot CLIP Radford et al.
+(2021), linear probing CLIP Radford et al.
+(2021), CoOp Zhou et al.
+(2022), VPT Jia et al.
+(2022) and DAPT Cho et al.
+(2023).
+CLIP+ZS CLIP+LP CoOp VPT DAPT Ours
+46.1 36.0 47.1 47.7 48.3 50.2
+Moreover, we evaluate the generalizability of our method in the domain generalization setting.
+Experiments are conducted on ImageNet-Sketch Wang et al.
+(2019) with ImageNet Russakovsky et al.
+(2015a) as the source dataset, as shown in the Table 3.
+From the experiment results, our method achieves superior performance with an accuracy of 50.2% on ImageNet-Sketch, attributed to the robustness of the T-distribution-based drift adapter.
+It further verifies the generalization ability of our model in the open world.
 3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
@@
 LT / Thp 36.2 ($\pm$3.53) 85.6 ($\pm$1.88) 101.3 ($\pm$1.25)
-Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training.
+Moreover, we verified at the feature level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training.
 As exhibited in Table 4, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/Thp.
@@
 Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter.
-It ensures our approach performs robust for OOD drift detection based on the spherical distribution robustly.
+It ensures our approach performs robustly for OOD drift detection.
 3.3 Ablation Experiments
 3.3.1 T-distributed Spherical Embedding in the Pre-training and Fine-tuning
-Firstly, we conduct ablation experiments to verify the improved performance of the T-distributed adapter in the stage of pre-training and fine-tuning, respectively.
-As demonstrated in Table 4, our proposed T-distributed adapter exhibits improvements in both the pre-training and fine-tuning stages of the vision language model.
-It is worth highlighting that the T-distribution adapter plays a more prominent role during the fine-tuning stage.
-We argue that it is due to different characteristics of the pre-training and fine-tuning.
-During fine-tuning, the model is directly involved in specific downstream tasks, and explicit category centers are present in the classifier.
-In contrast, pre-training primarily focuses on aligning image-text features, where the implicit information of categories is embedded.
-As a result, the T-distribution adapter’s impact is more pronounced in the fine-tuning stage compared to pre-training.
 Table 5: Ablation evaluation results with or without the T-distributed adapter in the pre-training or fine-tuning.
@@
 - 65.1
-69.6
-3.3.2 Various Concentration Parameter $\kappa$
-Furthermore, we conduct ablation experiments to examine the impact of different concentrations of the T-distributed adapter on the overall performance of the vision language model in Table 5.
-Four fixed degrees of concentration are involved, namely $\kappa=4$ 16, 64 and 128.
-The greater the degree of concentration, the greater the kurtosis of the T-distributed spherical metric.
-Besides, the concentration can also be utilized as a trainable parameter joining in the training, with the initial setting of 16.
-Table 5: Ablation evaluation results of various concentration parameter $\kappa$ on long-tailed classification task.
-"Training" denotes the $\kappa$ involved in the training as a parameter of the model with an initial setting of 16.
+69.4
+Firstly, we conduct ablation experiments to verify the improved performance of the T-distributed adapter in the stage of pre-training and fine-tuning, respectively.
+As demonstrated in Table 5, our proposed T-distributed adapter exhibits improvements in both the pre-training and fine-tuning stages of the vision language model.
+It is worth highlighting that the T-distribution adapter plays a more prominent role during the fine-tuning stage.
+We argue that it is due to different characteristics of the pre-training and fine-tuning.
+During fine-tuning, the model is directly involved in specific downstream tasks, and explicit category centers are present in the classifier.
+In contrast, pre-training primarily focuses on aligning image-text features, where the implicit information of categories is embedded.
+As a result, the T-distribution adapter’s impact is more pronounced in the fine-tuning stage compared to pre-training.
+3.3.2 Various Concentration $\kappa$ in T-Adapter
+Table 6: Ablation evaluation results of various concentrations of parameter $\kappa$ on the long-tailed classification task.
+”Training” denotes the $\kappa$ involved in the training as a parameter of the model with an initial setting of 16.
 The results are based on the ImageNet-LT with the Vit-base.
@@
 $\kappa$ 4 16 64 128 Training
-Acc 67.2 69.6 64.9 61.1 68.3
+Acc 67.2 69.4 64.9 61.1 68.3
+Furthermore, we conduct ablation experiments to examine the impact of concentrations of the T-distributed adapter on the overall performance of the VL models in Table 6.
+Four fixed degrees are involved, namely $\kappa$ = 4, 16, 64 and 128.
+The greater the degree of concentration, the greater the kurtosis of the Thp metric.
+Besides, the concentration can also be utilized as a trainable parameter joining in the training, with the initial setting of 16.
 In Table 6, setting the concentration parameter to $\kappa=16$ yields superior results.
-We argue that, a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model.
+We argue that a smaller concentration makes it challenging to effectively mitigate the biases introduced by tail drift and OOD drift in the vision language model.
 In terms of the bigger concentration, the model is hard to train due to the high kurtosis of the Thp metric.
 In the context of concentration as a trainable parameter with the initialization of 16, there is a slight reduction in model performance, accompanied by a marginal increase of the concentration parameter to 16.37.
-We assert that, the introduction of the new parameter increases the model’s complexity, thereby making the training process more challenging.
-Additionally, the low learning rate hinders the effective training of the concentration parameters, leading its less change compared with the initialization.
+We assert that the introduction of the new parameter increases the model’s complexity, thereby making the training process more challenging.
 4 Conclusions And Outlook
@@
 Additionally, we introduce a comprehensive set of multi-modal datasets (OpenMMlo) tailored to the long-tailed open world, which includes images, captions and related category annotations.
-We hope that our work will inspire future advancements in multi-modal large language models, specifically addressing the mitigation of biases originating from real-world data challenges, such as tailed drift and OOD drift.
-In future research, we will leverage concept drift theory to analyze the temporal, intensity, and regional aspects of the various drifts.
+Finally, we hope that our work will inspire future advancements in multi-modal large language models, specifically addressing the mitigation of biases originating from real-world data challenges, such as tailed drift and OOD drift.
+Acknowledgment
+The work was supported by the Australian Research Council (ARC) under Laureate project FL190100149.
+Appendix A Appendix
+A.1 Related Works
+A.1.1 Multi-modal Large Language Model
+Large Language Models (LLMs) have recently significantly impacted the field of natural language processing.
+Through alignment techniques such as supervised learning and reinforcement learning with human feedback, LLMs can effectively generalize to perform a wide range of tasks, even with limited training data.
+A remarkable application of LLM is ChatGPT, which presents an amazing ability to interact with humans.
+OpenAI’s ChatGPT and GPT4 are prime examples of the impact that AI can have, and there have been extensive open-source efforts to replicate their success, such as OPT Zhang et al.
+(2022), BLOOM Scao et al.
+(2022), PALM Chowdhery et al.
+(2022), LLaMA Touvron et al.
+(2023).
+Multi-modal large language models have further promoted the development of the vision-language models Radford et al.
+(2021); Li et al.
+(2022d); Alayrac et al.
+(2022); Li et al.
+(2023); Zhu et al.
+(2023); Liu et al.
+(2023a); Chen et al.
+(2023); Yang et al.
+(2024; 2025b).
+CLIP Radford et al.
+(2021) was introduced to separately extract features from the visual encoder and the text encoder, and combine them using contrastive learning.
+CLIP supports a variety of downstream tasks, including image retrieval, image classification tasks and especially zero-shot classification tasks.
+But, it cannot generate detailed captions based on images due to the lack of a text decoder.
+In contrast, our model primarily addresses the concept drift issue within multi-modal large language models, since an image-grounded text decoder is employed to generate text based on the images.
+Besides, CLIP requires a large-scale and high-quality WIT dataset to be driven, that contains 37.6 million entity image-text samples with 11.5 million unique images across 108 Wikipedia languages.
+Whereas, our method is validated under the extended ImageNet-LT, which consists of only 115.8K imbalanced images-text pairs.
+Building on CLIP, GLIP Li et al.
+(2022d) was developed to learn object-level, language-aware, and semantic-rich visual representations, unifying object detection and phrase grounding for pre-training.
+Different from the contrastive method, Flamingo Alayrac et al.
+(2022) aligned a pre-trained vision encoder and language model using gated cross-attention, demonstrating impressive few-shot learning capabilities.
+BLIP2 Li et al.
+(2023) was subsequently introduced, and it employed a Flan-T5 Chung et al.
+along with a Q-Former to effectively align visual features with the language model.
+MiniGPT4 Zhu et al.
+(2023), the most recent development in the field is the PaLM-E model, which features 562 billion parameters and is designed to integrate real-world continuous sensor modalities into an LLM, thereby establishing a connection between real-world perceptions and human languages.
+Based on Visual Fundamental Models like BLIP mentioned above, Visual ChatGPT adopts ChatGPT as the central component for interacting with users.
+It integrates multiple visual foundation models and utilizes prompt engineering, also known as Prompt Manager, to instruct ChatGPT about the usage, input-output format, and capabilities of each foundation model.
+This enables ChatGPT to determine how to invoke these models to fulfill the user’s requirements.
+Besides, GPT-4V(ision) OpenAI (2023) and GPT-4O(mni) have recently shown unprecedented ability in understanding and processing an arbitrary mix of input images and texts.
+A.1.2 Long-tailed Open World
+In vision tasks, significant efforts have been devoted to mitigating the challenges posed by the long-tailed open world.
+Two prominent research directions have emerged: long-tailed classification under open-world settings, exemplified by approaches like OLTR++ Liu et al.
+(2019; 2022b), LUNA Cai et al.
+(2022), DALC Wang et al.
+(2023), Open-sampling Wei et al.
+(2022) and TLC Li et al.
+(2022a), and OOD detection in long-tailed recognition, as seen in methods such as PASCL Wang et al.
+(2022), EAT Wei et al.
+(2024).
+OLTR++ Liu et al.
+(2019; 2022b) proposed an ensemble algorithm, consisting of dynamic meta-embedding to improve the recognition of tail categories and active learning for open categories detection.
+LUNA Cai et al.
+(2022) presented a distribution-sensitive loss to weigh more on the tail classes and a local-density-based metric to measure the novelty of OOD samples.
+DALC Wang et al.
+(2023) designed an active distribution optimization algorithm for clustering, querying and classification to balance the classification bias.
+Open-sampling Wei et al.
+(2022) rebalances class priors by sampling labels from a complementary distribution for each open-set instance, mitigating class imbalance.
+TLC Li et al.
+(2022a) utilizes the Dempster-Shafer Evidence Theory in a multi-expert framework for uncertainty estimation of tail and OOD samples.
+PASCL Wang et al.
+(2022) applied supervised contrastive learning to explicitly boost the model to distinguish between tail-class in-distribution samples and OOD samples.
+EAT Wei et al.
+(2024) introduces abstention classes for clear decision boundaries and augmenting tail classes with context-rich OOD data to focus on discriminative features.
+MCM Ming et al.
+(2022) pioneers the integration of vision language models into OOD detection, enabling zero-shot OOD by aligning visual features with text concepts through a proposed maximum concept matching approach.
+In addition, more and more VL methods have gained attention in the long-tail domain, such as LPT DONG et al.
+(2023), BALLAD Ma et al.
+(2021), Decoder Wang et al.
+(2024), VL-LTR Tian et al.
+(2022) and LIFT Shi et al.
+(2024).
+However, most of them pay attention to the fine-tuning of the vision language model under long-tailed scenarios.
+They directly use the pre-trained CLIP model, which is pre-trained using the high-quality and large-scale WIT dataset.
+In contrast, we are more concerned about the impact of long-tail open data on the whole model training from pre-training onwards, including pre-training and fine-tuning.
+Additionally, in the domain of the language model, Kandpal et al.
+(2023) corroborates that large language models (LLMs) also struggle to learn long-tailed knowledge.
+While larger models are better at absorbing long-tailed knowledge, they estimate that current models must be scaled by many orders of magnitude to reach competitive performance.
+Besides, Raunak et al.
+alleviates the long-tail problem in neural machine translation by quantifying token classification and sequence generation, and introduces an anti-focus loss that incorporates beam search inductive biases to better adapt model training to conditional text generation.
+A.1.3 Concept Drift
+In the review Lu et al.
+(2019), the algorithms related to concept drift are categorized into three groups: error rate-based, data distribution-based and multiple hypothesis-based.
+Our proposed algorithm belongs to the distribution-based concept drift detection and adaptation method.
+Distribution-based concept drift algorithms not only accurately detect drift through explicit distributions but also analyze the drift to identify its happening timing, location, and severity.
+Besides, RBM-IM Korycki & Krawczyk proposes a novel trainable concept drift detector based on Restricted Boltzmann Machine, to solve the concept drift in multi-class imbalanced data streams.
+Meanwhile, DDG-DA Li et al.
+initially trains a predictor to estimate future data distribution with concept drift, utilizes this information to create training samples, and subsequently trains models on the generated data.
+Furthermore, CALMID Liu et al.
+(2021) proposes a comprehensive active learning method for multiclass imbalanced streaming data with concept drift, including an ensemble classifier, a drift detector, and a variable threshold uncertainty strategy.
+Subsequently, DES-ICD Jiao et al.
+(2024) is a dynamic ensemble selection method for imbalanced data streams with concept drift.
+It considers the local performances of base classifiers and addresses class imbalance using a novel synthetic minority oversampling technique.
+Moreover, GOOD Gui et al.
+(2022) develops a graph OOD benchmark, which explicitly distinguishes between covariate and concept shifts and designs data splits that accurately capture these different shifts.
+Beyond that, ResilientCL Yang et al.
+(2025a) introduces a causal framework that integrates concept drift adaptation with structural causal modeling.
+By decoupling spurious correlations via causal graphs and enforcing counterfactual invariance, it addresses distributional biases in streaming training data.
+Besides, Liu et al.
+(2022a; 2023b; 2024) propose a multi-view uncertainty framework that addresses concept drift across heterogeneous data streams through set-valued prediction generation, effectively consolidating probabilistic outputs into deterministic categorical representations.
+Remark A.1.
+Differences: Concept Drift vs.
+Data Drift (Covariate Drift) Data drift entails changes solely in the distribution of inputs $P(x)$, while concept drift involves alterations in both input and output distributions, i.e., $P(x)$ and $P(y)$, leading to changes in the decision boundary.
+Furthermore, data drift predominantly stems from internal factors like data collection and processing, whereas concept drift typically arises from external factors, reflecting real-world changes.
+A.1.4 Hyperspherical Distribution Modelling
+The Bayesian estimation of the vMF mixture model with variational inference is addressed in Taghia et al.
+.
+The learning task in VI consists of the optimization of the variational posterior distribution.
+Besides, a deep metric learning model for image classification and retrieval is presented in Zhe et al.
+, which utilizes the vMF distribution to define the loss function and introduces an effective alternative learning algorithm by updating class centers.
+The model captures global information in the embedding space and approximates the class distribution during training, leading to improved performance in image tasks.
+Kobayashi extends the vMF distribution to regularize the intra-class feature distribution for imbalanced, small-scale and noisy data.
+Yang et al.
+(2023) focus on using hyperspherical embedding to alleviate the crowding problem arisen by the imbalanced data.
+Ming et al.
+(2023) utilizes hyperspherical embeddings for OOD detection in representation learning, consisting of two losses, a dispersion loss to increase angular distances between different class prototypes, and a compactness loss to ensure samples are closer to their respective class prototypes.
+Besides, H-SRDC Tang et al.
+enhances intra-class compactness by combining target data clustering with a domain-shared classifier and cluster centroid learning, enhancing deep clustering by minimizing Kullback-Leibler divergence between network predictions and an auxiliary distribution.
+A.2 The T-distributed Distribution on Hypersphere
+A.2.1 Directional statistics
+Directional statistics primarily focus on the distribution of eigenvector angles, while neglecting the impact of eigenvector module lengths.
+Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set.
+A key idea in directional distribution is the tangent-normal decomposition.
+Any unit vector $x$ can be decomposed as:
+$x=t\mu+(1-t^{2})^{\frac{1}{2}}v,t\in[-1,1],$ (9)
+with $v\in\mathbb{S}^{d-2}$ a tangent to $\mathbb{S}^{d-1}$ at $\mu$ Mardia & Jupp (2000); De Cao & Aziz (2020), where $v$ and $t$ are independent and $v$ is uniform on $\mathbb{S}^{d-2}$.
+Thus, the intersection of $\mathbb{S}^{d-1}$ with the hyperplane through $t\mu$ and normal to $\mu$ is a $(d-2)$-dimensional sphere of radius $\sqrt{1-t^{2}}$, that $t$ has density as following:
+$p_{T}(t;d)\propto(1-t^{2})^{\frac{d-3}{2}},t\in[-1,1].$ (10)
+Therefore, through the marginal density $p_{T}$ and $p_{v}$, we can estimate the density of the entire spherical distribution.
+One prominent instance is the von Mises-Fisher distribution (vMF) Banerjee et al.
+(2005), which can be interpreted as a probability distribution over the cosine similarity between a unit vector $x$ and a fixed mean direction $\mu$, following the density:
+$p_{X}(x;\mu,\kappa)\propto\exp{(\kappa\mu^{T}x)},$ (11)
+where $\kappa\geqslant 0$ denotes the concentration and $\exp$ represents the exponential function. Therefore, combined with the Eq. 9 and Eq. 10, the density of vMF is:
+$p(x)=C_{X}(\kappa,d)^{-1}\exp{(\kappa\mu^{T}x)},\quad x\sim\text{vMF}(\mu,\kappa)$ (12)
+$C_{X}(\kappa,d)=\frac{(2\pi)^{d/2}I_{d/2-1}(\kappa)}{\kappa^{d/2-1}},$
+where $I_{m}$ denotes the modified Bessel function of the first kind at order $m$.
+A.2.2 Derivation of the T-distributed Distribution on Hypersphere
+Given the unit feature vector $X_{ij}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:||x||_{2}=1\}$ denotes the $(d-1)$-dimensional hyperspherical set.
+The proposed T-distribution metric on hypersphere (Thp) follows the density:
+$p_{X}(x)\propto\frac{2}{\kappa(1-\mu^{T}x)},$ (13)
+where $x\in\mathbb{S}^{d-1}$, direction $\mu in\mathbb{S}^{d-1}$ and concentration $\kappa\in\mathbb{R}_{\geq 0}$.
+Let $T$ bet a random variable that denotes the dot-product $t=\mu^{T}x$, then $T=2Z-1$, with $Z\sim\text{Beta}(\alpha,\beta)$, where $\alpha=\frac{d-1}{2}$ and $\frac{d-3}{2}$.
+Proof.
+Given Eq. 10, the marginal distribution of the dot-product $t$ is
+$t\propto\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}.$ (14)
+So, its normalizer is:
+$N_{T}(\kappa,d)$ $=\int_{\mathbb{S}^{d-1}}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}\mathrm{d}t$ (15)
+$=\int_{-1}^{1}\frac{1}{\kappa(1-t)}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-3}{2}}\mathrm{d}t$
+$=\frac{1}{\kappa}\int_{-1}^{1}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}\mathrm{d}t.$
+Given the useful integral function:
+$\int(1+x)^{a}(1-x)^{b}\mathrm{d}x=2^{a+b+1}B_{\frac{x+1}{2}}(a+1,b+1)+C.$ (16)
+So, its normalizer is:
+$N_{T}(\kappa,d)$ $=\frac{1}{\kappa}2^{d-3}(B_{1}(\frac{d-1}{2},\frac{d-3}{2})-B_{0}(\frac{d-1}{2},\frac{d-3}{2}))$ (17)
+$=\frac{1}{\kappa}2^{d-3}B(\frac{d-1}{2},\frac{d-3}{2}).$
+The Beta function:
+$B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}.$ (18)
+So, the normalizer is
+$N_{T}(\kappa,d)=\frac{1}{\kappa}2^{\alpha+\beta-1}\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)},$ (19)
+where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. It follows that the probability density function of the marginal distribution of the dot product is,
+$p_{T}(t;\kappa,d)$ $=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa(1-t)}(1-t^{2})^{\frac{d-3}{2}}$ (20)
+$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(1+t)^{\frac{d-3}{2}}(1-t)^{\frac{d-5}{2}}$
+$=N_{T}(\kappa,d)^{-1}\frac{2}{\kappa}(2z)^{\frac{d-1}{2}-1}(2-2z)^{\frac{d-3}{2}-1}$
+$=\frac{2}{\kappa}B(\alpha,\beta)^{-1}z^{\alpha-1}(1-z)^{\beta-1},$
+where, $\alpha=\frac{d-1}{2}$ and $\beta=\frac{d-3}{2}$. ∎
+Due to the surface area of the hyper-sphere $\mathbb{S}^{d-1}$ is:
+$A_{d-1}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}.$ (21)
+The T-distributed spherical distribution is expressed via the tangent normal decomposition as a joint distribution between $T\sim p_{T}{t;\kappa,d}$ and $V\sim\mathcal{U}(\mathbb{S}^{d-2})$.
+Since $T\perp\!\!\!\perp V$, the Thp normalizer $N_{x}(p,k)$ is the product of the normalizer of $p_{T}(t;\kappa,d)$ and the uniform distribution on $\mathbb{S}^{d-2}$ is:
+$N_{X}(\kappa,d)$ $=N_{T}(\kappa,d)\cdot A_{d-2}$ (22)
+$=2^{\alpha+\beta-1}B(\alpha,\beta)\frac{2\pi^{\beta}}{\kappa\Gamma(\beta)}$
+$=\frac{2^{\alpha+\beta}\pi^{\beta}}{\kappa}\frac{\Gamma(\alpha)}{\Gamma(\alpha+\beta)}.$
+Thus,
+$p_{X}(x;\mu,\kappa)=N_{X}(\kappa,d)^{-1}\frac{2}{\kappa(1-\mu^{T}x)}.$ (23)
+A.3 Implementation Details
+For our language-guided image tokenizer, we leverage the strengths of both BERT Devlin et al. (2019b) and ViT as our text encoder, text decoder and visual encoder, respectively.
+We employ ViT-Bae as our visual encoder, which consists of 12 transformer encoder layers and an FFN intermediate size of 3,072.
+The input image size is set to $384\times 384$, with a patch size of $16\times 16$.
+The hidden dimensions of the ViT-Base are 768, with 12 attention heads.
+And, the number of parameters is about 86 million.
+Besides, we also use ResNeXt-50 to perform ablation experiments.
+In addition, ResNeXt-50 has 16 residual blocks with 50 layers.
+Each block has 3 convolutional layers with the kernel size of $3\times 3$, the stride of 1 and the padding of 1.
+The batch normalization and max pooling are utilized to connect the convolutional layers.
+The classification head hidden dimensions are 2,048.
+Additionally, BERT as the language model in our vision-language model, has 12 transformer layers with 768 hidden dimensions and 3,078 intermediate dimensions.
+The number of attention heads is 12, with the input sequence length of 512.
+It has approximately 110 million parameters.
+(a) Sample in Training Set
+(b) Sample in Test Set
+(c) Sample in Open Set
+Figure 4: Samples of OpenMMlo in training set, test set and open set.
+In terms of the pre-training progress, the hyperparameters are presented in Table 7.
+We utilize the AdamW optimizer, which is configured with a cosine annealing schedule as the learning policy.
+The initial learning rate is set to $2\times 10^{-5}$, and the AdamW optimizer is employed with hyperparameters $\beta=(0.9,0.98)$.
+Additionally, we set the weight decay to 0.05 and the dropout rate to 0.1.
+During the first 1,000 warm-up steps, the learning rate increases to $2\times 10^{-5}$, and subsequently decays to $10^{-7}$.
+Unless otherwise specified, the pre-training of our vision language model consists of 800,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs.
+And the pre-training experiments are conducted in the manner of different stages, namely gradual drifts with long-tailed data and sudden drifts with OOD data.
+It is mainly to compare with different methods with the same setup.
+Table 7: The training hyperparameters of our vision language model.
+Pre-training
+Training Steps 400,000
+Warmup Steps 1,000
+Optimizer AdamW
+Learning Rate 1e-4
+Learning Rate Decay Cosine
+Adam $\beta$ (0.9, 0.98)
+Weight Decay 0.05
+Batch Size 50
+Fine-tuning
+Training Steps 18,000
+Warmup Steps 0
+Optimizer AdamW
+Learning Rate 2e-5
+Learning Rate Decay Cosine
+Adam $\beta$ (0.9, 0.98)
+Weight Decay 0.05
+Batch Size 400
+While in the fine-tuning on the downstream task of classification, the initial learning rate is reduced to $10^{-6}$ without the warmup.
+The visual encoder and text decoder are frozen out of the training.
+Thus, the batch size can be increased to 400.
+The fine-tuning consists of 18,000 steps, executed on $2\times 2$ NVIDIA A100 GPUs.
+Other training parameters are the same as the pre-training.
+Besides, under the only fine-tuning settings, the image encoder and the text encoder are frozen with the CLIP pre-trained parameters, while the image-grounded text decoder is trained during the fine-tuning.
+When evaluating the performance of our VL model under the long-tailed open world, we use the top-1 accuracy metric on the downstream classification task.
+In particular, the categories are split into three groups: many-shot (with more than 100 training samples), medium-shot (with 20-100 training samples), and few-shot (with fewer than 20 training samples).
+The Top-1 accuracies are computed for each group to evaluate the performance of mitigating the bias introduced by the long-tail distribution, respectively.
+Furthermore, in order to assess the capability of detecting the OOD drift, we employ two metrics: FPR95 which measures the false positive rate of OOD samples when the true positive rate of ID samples reaches 95%, and AUROC providing the area under the receiver operating characteristic curve.
+Besides, cosine distance is exploited to measure the distances between features and centers in the feature space of the VL model.
+A.4 Building Multi-modal Long-Tailed OOD Datasets Group OpenMMlo
+Figure 4 showcases the samples utilized for training and validation in our study.
+To intuitively verify the impact of long-tail open-world scenarios on multi-modal large language models, we employ classification as our downstream task.
+When matching images and texts, we strategically mask words that are directly related to category names.
+This approach ensures the accuracy and reliability of our experimental results.
+As depicted in Figure 4, comprehensive descriptions of the image are provided through long-form text, encompassing details such as size, position, color, relationships, and other relevant information about the objects present in the image.
+This ensures a detailed and information-rich depiction of the visual content.
+We have publicly released the datasets used for training and validation, as well as the original unmasked datasets.